### v3

# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

Abstract
The optimization foundations of deep linear networks have recently received significant attention. However, due to their inherent non-convexity and hierarchical structure, analyzing the loss functions of deep linear networks remains a challenging task. In this work, we study the local geometric landscape of the regularized squared loss of deep linear networks around each critical point. Specifically, we derive a closed-form characterization of the critical point set and establish an error bound for the regularized loss under mild conditions on network width and regularization parameters. Notably, this error bound quantifies the distance from a point to the critical point set in terms of the current gradient norm, which can be used to derive linear convergence of first-order methods. To support our theoretical findings, we conduct numerical experiments and demonstrate that gradient descent converges linearly to a critical point when optimizing the regularized loss of deep linear networks.
Keywords: Deep linear networks, critical points, error bounds, linear convergence
1 Introduction
Deep learning has achieved great empirical success in various fields, including computer vision [17], natural language processing [41], and healthcare [13]. Optimization is a key component of deep learning, playing a pivotal role in formulating learning objectives and training neural networks. In general, optimization problems arising in deep learning are highly non-convex and difficult to analyze due to the inherent nonlinearity and hierarchical structures of deep neural networks. Even in the context of linear neural networks, which are the most basic form of neural networks, our theoretical understanding remains far from complete and systematic. In this work, we study the following problem based on deep linear networks:
$\min_{\bm{W}}\ F(\bm{W}):=\left\|\bm{W}_{L}\cdots\bm{W}_{1}-\bm{Y}\right\|_{F}^{2}+\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}\|_{F}^{2},$ (1)
where $L\geq 2$ denotes the number of layers, $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$ denotes the data input, $\bm{W}_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ denotes the $l$-th weight matrix, $\bm{W}=\{\bm{W}_{l}\}_{l=1}^{L}$ denotes the collection of all weight matrices, and $\lambda_{l}>0$ for all $l$ are regularization parameters. Notably, such a problem captures a wide range of deep learning problems, including deep matrix factorization [2, 8, 40], neural collapse [15, 50], and low-rank adaptation of learning models [19, 45].
In the literature, a large amount of work has been dedicated to studying the following problem:
$\min_{\bm{W}}\left\|\bm{W}_{L}\cdots\bm{W}_{1}\bm{X}-\bm{Z}\right\|_{F}^{2},$
where $(\bm{X},\bm{Z})\in\mathbb{R}^{d_{0}\times N}\times\mathbb{R}^{d_{L}\times N}$ denotes the data input. A common assumption in many theoretical studies (see, e.g., [1, 5, 9, 44]) is that the data is whitened, i.e., $\bm{X}\bm{X}^{T}=\bm{I}_{d}$. Under this assumption, one can verify that $\min_{\bm{W}}\left\|\bm{W}_{L}\cdots\bm{W}_{1}\bm{X}-\bm{Z}\right\|_{F}^{2}$ is equivalent to $\min_{\bm{W}}\left\|\bm{W}_{L}\cdots\bm{W}_{1}-\bm{Z}\bm{X}^{T}\right\|_{F}^{2}$. Obviously, this formulation is an unregularized counterpart of Problem (1). We should point out that regularization is commonly used during the training of neural networks to prevent overfitting, improve generalization, and accelerate convergence [25, 48]. However, the existing theoretical studies mainly focus on unregularized deep neural networks and cannot be directly applied to Problem (1), even when the data is whitened; see, e.g., [1, 4, 5, 7, 20, 32, 44].
In practice, (stochastic) gradient descent (GD) and its variants are among the most widely used first-order methods for deep learning [24]. Over the past few years, substantial progress has been made in studying the convergence behavior of GD for optimizing deep linear networks, particularly in the context of the unregularized problem $\min_{\bm{W}}\left\|\bm{W}_{L}\cdots\bm{W}_{1}-\bm{Y}\right\|_{F}^{2}$. For example, Bartlett et al. [5] showed that GD with the identity initialization converges to an $\epsilon$-optimal solution within a polynomial number of iterations. Later, Arora et al. [1] further improved the convergence result, showing that GD converges linearly to a global optimum when $\min\{d_{1},\dots,d_{L-1}\}\geq\min\{d_{0},d_{L}\}$, the initial weight matrices are approximately balanced, and the initial loss is smaller than a threshold. Meanwhile, Wu et al. [44] showed that gradient descent with zero-asymmetric initialization avoids saddle points and converges to an $\epsilon$-optimal solution in $O(L^{3}\log(1/\epsilon))$ iterations. Other works also proved similar global convergence and convergence rate results of GD under different assumptions; see, e.g., [4, 7, 20, 32, 44]. Despite these inspiring results, the existing analyses suffer from some notable limitations. First, they are highly specific to a particular problem, relying on the analysis of gradient dynamics under carefully designed weight initialization schemes. In addition, the existing analyses only apply to analyze the convergence to global optimal solutions. However, Problem (1) and its unregularized counterpart may have other local solutions, to which first-order methods, such as GD, are likely to converge. To the best of our knowledge, the convergence behavior of first-order methods when they approach a critical point—whether a global minimum, a local minimum, or even a saddle point—remains an open question in the literature. To sum up, it remains an unsolved challenge in the literature to develop a unified framework to analyze the convergence behavior of first-order methods when solving Problem (1).
To address the above challenge, a promising approach is to study the local geometric structure of Problem (1) associated with its objective function, such as the error bound condition [29, 36, 51], the Polyak-Łojasiewicz (PŁ) inequality [34], and quadratic growth [10]. In this work, we mainly focus on the error bound condition associated with Problem (1). Formally, let $\mathcal{W}_{F}$ denote the set of critical points of Problem (1). We say that it possesses an error bound for $\mathcal{W}_{F}$ if there exist constants $\epsilon,\kappa>0$ such that for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\epsilon$, we have
$\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\kappa\|\nabla F(\bm{W})\|_{F},$ (2)
where $\mathrm{dist}(\bm{W},\mathcal{W}_{F}):=\min\{\|\bm{W}-\bm{V}\|_{F}:\bm{V}\in\mathcal{W}_{F}\}$ denotes the distance from $\bm{W}$ to the critical point set $\mathcal{W}_{F}$. Intuitively, the error bound inequality (2) requires the distance from a point to the set of critical points to be bounded by its gradient norm. A unified framework leveraging this condition, along with some algorithm-dependent conditions, has been widely studied to analyze linear convergence of first-order methods [10, 51] or superlinear convergence of second-order methods [47] in convex optimization. Luo and Tseng [31] showed that even if the objective function is non-convex but smooth and satisfies the error bound, (projected) GD converges linearly to the solution set. Moreover, Li and Pong [28], Liao et al. [29], Rebjock and Boumal [36] demonstrated that the error bound is equivalent to other regularity conditions, such as the PŁ inequality [34] and quadratic growth [10] under mild conditions. These regularity conditions are widely used in the literature to prove linear convergence of GD for optimizing non-convex problems [34, 31, 3, 14]. Notably, this type of convergence analysis framework is not limited to studying convergence to global optimal solutions. Instead, it can also be applied to local minima or even saddle points, as long as the error bound holds for the targeted solution sets. This makes such a framework particularly effective for analyzing non-convex optimization problems, whose landscape often contains various types of critical points.
However, powerful as this approach may seem, proving error bounds for specific problems, particularly non-convex ones, is technically difficult, as it often requires a delicate analysis [51, 21, 22, 30, 43, 33]. As far as we know, proving the error bound or other regularity conditions for deep neural networks remains relatively underexplored. Recently, Wang et al. [42] made progress in this direction by establishing the error bound for the set of global optimal solutions to a special instance of Problem (1): $\min\|\bm{W}_{2}\bm{W}_{1}-\bm{Y}\|_{F}^{2}+\lambda_{1}\|\bm{W}_{1}\|_{F}^{2}+\lambda_{2}\|\bm{W}_{2}\|_{F}^{2}$, where $\lambda_{1},\lambda_{2}>0$ are regularization parameters and $\bm{Y}$ is a membership matrix. Moreover, Zhang et al. [49] and Tao et al. [39] extended the analysis to the setting of general data input $\bm{Y}$. However, these analyses are limited to global optimal solutions of 2-layer linear networks. Extending these results to deeper networks and critical points remains an open problem.
In view of the above discussion, our goal in this paper is to establish an error bound for the set of critical points of Problem (1). Our main contributions are twofold. First, we explicitly characterize the critical point set of Problem (1) for general data $\bm{Y}$ (see 1), despite the non-convexity and hierarchical structure of the problem. This characterization serves as a foundation for establishing the error bound for Problem (1). Second, leveraging this explicit characterization, we show that all critical points of Problem (1) satisfy the error bound (see 2) under mild conditions on the network width and regularization parameters. More specifically, we identify the sufficient and necessary conditions on the relationship between the regularization parameters and the spectrum of the input data that ensure the error bound holds. Such a result is significant, as it expands the currently limited repertoire of non-convex problems for which the local loss geometry is well understood. Moreover, it is important to note that our work develops new techniques in the process of establishing the error bound to handle the repeated singular values in $\bm{Y}$ and the complicated structure of the critical point set. These techniques could be of independent interest. The established error bound can be used to establish other regularity conditions, such as the PŁ inequality and quadratic growth. Building on the error bound of Problem (1), we demonstrate that first-order methods can achieve linear convergence to a critical point of Problem (1), provided that they satisfy certain algorithm-dependent properties. Finally, we conduct numerical experiments in deep linear networks and in more general settings and observe that gradient descent converges linearly to critical points. These results strongly support our theoretical findings.
Notation. Given a positive integer $n$, let $n]:=\{1,\dots,n\}$. Given a set of numbers $\{\lambda_{1},\dots,\lambda_{n}\}$, we define $\lambda_{\min}:=\min\{\lambda_{i}:i\in[n]\}$ and $\lambda_{\max}:=\max\{\lambda_{i}:i\in[n]\}$. Given a vector $\bm{a}$, let $\|\bm{a}\|$ denote its Euclidean norm, $\|\bm{a}\|_{0}$ its $\ell_{0}$ norm (i.e., number of non-zero entries), $a_{i}$ the $i$-th entry, and $\mathrm{diag}(\bm{a})$ the diagonal matrix with $\bm{a}$ on the diagonal. Given a matrix $\bm{A}$, let $\|\bm{A}\|$ denote the spectral norm, $\|\bm{A}\|_{F}$ the Frobenius norm, $a_{ij}$ the $(i,j)$-th element, and $\sigma_{i}(\bm{A})$ the $i$-th largest singular value. We use $\bm{0}_{m\times n}$ to denote an all-zero matrix of size $m\times n$, $\bm{0}_{m}$ to denote all-zero vector of dimension $m$, and simply write $\bm{0}$ when the dimension can be inferred from the context. We denote the $i$-th standard basis vector by $\bm{e}_{i}$. We simply write $\nabla_{\bm{W}_{l}}F(\bm{W})$ as $\nabla_{l}F(\bm{W})$.
We use $\mathcal{O}^{n\times d}$ to denote the set of all $n\times d$ matrices that have orthonormal columns (in particular, $\mathcal{O}^{n}$ for $n\times n$ orthogonal matrices); $\mathcal{P}^{n}$ to denote the set of all $n\times n$ permutation matrices; $\mathrm{BlkD}(\bm{X}_{1},\dots,\bm{X}_{n})$ to denote the block diagonal matrix whose diagonal blocks are $\bm{X}_{1},\dots,\bm{X}_{n}$. Given a matrix $\bm{X}\in\mathbb{R}^{m\times n}$ and a non-empty closed set $\mathcal{X}\subseteq\mathbb{R}^{m\times n}$, the Euclidean distance from $\bm{X}$ to $\cal X$ is defined as $\mathrm{dist}(\bm{X},\mathcal{X})=\min_{\bm{Z}\in\mathcal{X}}\|\bm{X}-\bm{Z}\|_{F}$; the distance between $\mathcal{X}$ and a closed set $\mathcal{Y}\neq\emptyset$ is defined as $\mathrm{dist}(\mathcal{X},\mathcal{Y})=\min_{\bm{X}\in\mathcal{X},\bm{Y}\in\mathcal{Y}}\|\bm{X}-\bm{Y}\|_{F}$. Let $\pi:[d]\rightarrow[d]$ denote a permutation of the elements in $[d]$ with the corresponding permuation matrix $\bm{\Pi}\in\mathcal{P}^{d}$ satisfying $\bm{\Pi}(i,j)=1$ if $j=\pi(i)$ and $\bm{\Pi}(i,j)=0$ otherwise for each $i\in[d]$. Given weight matrices $\{\bm{W}_{l}\}_{l=1}^{L}$ with $\bm{W}_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ for each $l\in[L]$, let $\bm{W}_{i:1}:=\bm{W}_{i}\bm{W}_{i-1}\cdots\bm{W}_{1}$ for each $i=2,\dots,L$ and $\bm{W}_{L:i}:=\bm{W}_{L}\bm{W}_{L-1}\cdots\bm{W}_{i}$ for each $i=1,\dots,L-1$. In particular, we define $\bm{W}_{0:1}:=\bm{I}$ and $\bm{W}_{L:L+1}:=\bm{I}$. For all $i\geq j+1$, let $\bm{W}_{i:j}=\bm{W}_{i}\bm{W}_{i-1}\cdots\bm{W}_{j}$ and $\bm{W}_{i:j}=\bm{W}_{i}$ when $i=j$. All other notations are standard.
Organization. In Section 2, we present the main results of this paper. In Section 3, we present the proofs of the main results. We report experimental results in Section 4 and conclude the paper in Section 5.
2 Main Results
To begin, we define the set of critical points of Problem (1) as follows:
$\mathcal{W}_{F}:=\left\{\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L}):\nabla F(\bm{W})=\bm{0}\right\}.$
Let $\bm{Y}=\bm{U}_{Y}\bm{\Sigma}_{Y}\bm{V}_{Y}^{T}$ be an SVD of $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$, where $\bm{U}_{Y}\in\mathcal{O}^{d_{L}}$ and $\bm{V}_{Y}\in\mathcal{O}^{d_{0}}$. In addition,
$\bm{\Sigma}_{Y}=\mathrm{BlkD}\left(\mathrm{diag}(y_{1},y_{2},\dots,y_{d_{\min}}),\bm{0}_{(d_{L}-d_{\min})\times(d_{0}-d_{\min})}\right)\in\mathbb{R}^{d_{L}\times d_{0}},$ (3)
where $d_{\min}:=\min\{d_{0},d_{L}\}$ and $y_{1}\geq y_{2}\geq\dots\geq y_{d_{\rm min}}\geq 0$ are singular values. Here, if $d_{\min}=d_{0}$ (resp., $d_{\min}=d_{L}$), the lower diagonal block vanishes, leaving a $\bm{0}$ block of the appropriate size at the bottom (resp., on the right). Specifically, if $d_{\min}=d_{0}$, $\bm{\Sigma}_{Y}=\begin{bmatrix}\mathrm{diag}(y_{1},y_{2},\dots,y_{d_{\min}})\\ \bm{0}_{(d_{L}-d_{\min})\times d_{0}}\end{bmatrix}$; otherwise, $\bm{\Sigma}_{Y}=\begin{bmatrix}\mathrm{diag}(y_{1},y_{2},\dots,y_{d_{\min}})&\bm{0}_{d_{L}\times(d_{0}-d_{\min})}\end{bmatrix}$. To handle the repeated singular values, we introduce the following setup. Let $r_{Y}:=\mathrm{rank}(\bm{Y})$ and $p_{Y}$ denote the number of distinct positive singular values of $\bm{Y}$. That is, there exist indices $s_{0},s_{1},\dots,s_{p_{Y}}$ such that $0=s_{0}<s_{1}<\dots<s_{p_{Y}}=r_{Y}$ and
$\textstyle y_{s_{0}+1}=\dots=y_{s_{1}}>y_{s_{1}+1}=\dots=y_{s_{2}}>\dots>y_{s_{p_{Y}}+1}=\cdots=y_{d_{\text{min}}}=0.$ (4)
Note that if $r_{Y}=d_{\min}$, then $y_{s_{p_{Y}}}=y_{d_{\min}}>0$. Let $h_{i}:=s_{i}-s_{i-1}$ denote the multiplicity of the $i$-th largest positive singular value for each $i\in[p_{Y}]$. Throughout the paper, we consistently use the above notation.
2.1 Characterization of the Critical Point Set
To begin, we introduce the following assumption on the widths of network layers.
Assumption 1.
The network widths satisfy $\min\{d_{1},\dots,d_{L-1}\}\geq\min\{d_{0},d_{L}\}$.
This assumption ensures that the width of each hidden layer is no less than that of the input or output layer. Notably, it aligns with common practices in deep learning, where hidden layers are typically designed to be wide to have sufficient capacity to capture complex data representations [16, 35]. In addition, this assumption is widely used in the literature to analyze the convergence behavior of GD for optimizing deep networks; see, e.g., [1, 11, 12, 26]. Then we derive the closed-form expression of the critical point set $\mathcal{W}_{F}$ as follows:
Theorem 1 (Characterization of critical points).
Suppose that Assumption 1 holds. It holds that $\bm{W}\in\mathcal{W}_{F}$ if and only if
$\bm{W}_{1}=\bm{Q}_{2}\bm{\Sigma}_{1}\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right)\bm{V}_{Y}^{T},$ (5a)
$\bm{W}_{l}=\bm{Q}_{l+1}\bm{\Sigma}_{l}\bm{Q}_{l}^{T},\ l=2,\dots,L-1,$ (5b)
$\bm{W}_{L}=\bm{U}_{Y}\mathrm{BlkD}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{L}\bm{Q}_{L}^{T},$ (5c)
where $\bm{Q}_{l}\in\mathcal{O}^{d_{l-1}}$ for all $l=2,\dots,L$, $\bm{O}_{i}\in\mathcal{O}^{h_{i}}$ for each $i\in[p_{Y}]$, $\bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}}$, $\widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}}$, and $\bm{\Sigma}_{l}=\mathrm{BlkD}\left(\mathrm{diag}(\bm{\sigma})/\sqrt{\lambda_{l}},\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}}$ for each $l\in[L]$ with $\bm{\sigma}\in\mathbb{R}^{r_{Y}}$ satisfying
$\textstyle\sigma_{i}^{2L-1}-\sqrt{\lambda}y_{i}\sigma_{i}^{L-1}+\lambda\sigma_{i}=0,\ \sigma_{i}\geq 0,\ \forall i\in[r_{Y}],\ \text{where}\ \lambda:=\prod_{l=1}^{L}\lambda_{l}.$ (6)
We remark that this theorem is a special case of [6, Theorem 2.1], which provides a closed-form characterization of the critical point set of Problem (1) without any conditions on the network widths. In contrast, by leveraging Assumption 1, we obtain a simplified and more explicit description of the critical point set in this theorem. This characterization serves as a cornerstone for establishing the error bound for Problem (1). In general, these critical points are not isolated points but form a union of connected sets (see 1). This brings significant difficulty in computing the distance $\mathrm{dist}(\bm{W},\mathcal{W}_{F})$ from a point $\bm{W}$ to the critical point set, as this distance does not admit a closed-form solution due to the complex form of $\mathcal{W}_{F}$. One of our key contributions is to develop a method that overcomes this difficulty.
2.2 Error Bound for Deep Linear Networks
Before we proceed, it is worth noting that there are some degenerate cases under which the error bound does not hold. For example, when $L=2$, we study
$\min_{\bm{W}_{1},\bm{W}_{2}\in\mathbb{R}^{d\times d}}F(\bm{W})=\|\bm{W}_{2}\bm{W}_{1}-\lambda\bm{I}_{d}\|_{F}^{2}+\lambda\left(\|\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{2}\|_{F}^{2}\right).$
Applying 1 to the above problem yields $\mathcal{W}_{F}=\{(\bm{0},\bm{0})\}$. Now, consider $\bm{W}_{1}=\bm{W}_{2}=x\bm{I}_{d}$. We compute
$\nabla_{1}F(\bm{W})=2\bm{W}_{2}^{T}\left(\bm{W}_{2}\bm{W}_{1}-\lambda\bm{I}_{d}\right)+2\lambda\bm{W}_{1}=2x^{3}\bm{I}_{d},$
$\nabla_{2}F(\bm{W})=2\left(\bm{W}_{2}\bm{W}_{1}-\lambda\bm{I}_{d}\right)\bm{W}_{1}^{T}+2\lambda\bm{W}_{2}=2x^{3}\bm{I}_{d}.$
Therefore, we have $\|\nabla F(\bm{W})\|_{F}^{2}=8dx^{6}$. Moreover, we compute $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F})=\|\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{2}\|_{F}^{2}=2dx^{2}$. Consequently, we have $\sqrt{d}\|\nabla F(\bm{W})\|_{F}=\mathrm{dist}^{3}(\bm{W},\mathcal{W})$. Obviously, the error bound (2) does not hold as $x\to 0$. To avoid such degenerate cases, we impose the following conditions on the relationship between the regularization parameters and the data matrix $\bm{Y}$.
Assumption 2.
It holds for $L=2$ that
$\lambda:=\lambda_{1}\lambda_{2}\neq y_{i}^{2},\ \forall i\in[r_{Y}].$ (7)
For all $L\geq 3$, it holds that
$\lambda:=\lambda_{1}\cdots\lambda_{L}\neq y_{i}^{2(L-1)}\left(\left(\frac{L-2}{L}\right)^{\frac{L}{2(L-1)}}+\left(\frac{L}{L-2}\right)^{\frac{L-2}{2(L-1)}}\right)^{-2(L-1)},\ \forall i\in[r_{Y}].$ (8)
Notably, this assumption provides necessary and sufficient conditions under which the error bound holds for all critical points of Problem (1). Specifically, if Assumption 2 does not hold, the error bound fails to hold (see Appendix A). Conversely, if Assumption 2 holds, the error bound holds as follows:
Theorem 2 (Error bound).
Suppose that Assumptions 1 and 2 hold. There exist constants $\epsilon_{1},\kappa_{1}>0$ such that for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\epsilon_{1}$, we have
$\mathrm{dist}\left(\bm{W},\mathcal{W}_{F}\right)\leq\kappa_{1}\|\nabla F(\bm{W})\|_{F}.$ (9)
This theorem is significant as it shows that the error bound holds at any critical point of Problem (1) under Assumptions 1 and 2. Importantly, the constants $\kappa_{1},\epsilon_{1}$ can be explicitly derived through our proofs (see 2, 3, and (125) and (126) in the proof of 2). Now, we discuss some implications of our result. First, the established error bound can be used to derive other regularity conditions, such as PŁ inequality (see (10)) and quadratic growth (see (11)), for analyzing the convergence behavior of first-order methods. This result is well known; see, e.g., [36].
Corollary 1.
Suppose that $\bm{W}^{*}$ is a critical point and (9) holds for all $\bm{W}$ satisfying $\|\bm{W}-\bm{W}^{*}\|_{F}\leq\epsilon_{1}$.
(PŁ inequality) There exists a constant $\mu_{1}>0$ such that
$\|\nabla F(\bm{W})\|_{F}^{2}\geq\mu_{1}(F(\bm{W})-F(\bm{W}^{*})).$ (10)
(ii) (Quadratic growth) If $\bm{W}^{*}$ is a local minimizer, there exists a constant $\mu_{2}>0$ such that
$\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F})\leq\mu_{2}\left(F(\bm{W})-F(\bm{W}^{*})\right).$ (11)
Second, by combining the error bound with other algorithm-dependent conditions, such as sufficient decrease, cost-to-go estimate, and safeguard, we can establish local linear convergence of first-order methods for solving Problem (1) (see Appendix C). In contrast to algorithm-specific convergence rate analyses in [4, 7, 20, 32, 44], the error-bound-based framework provides a unified analytical approach. In particular, it applies not only to GD but also to a broad class of first-order methods capable of optimizing Problem (1). Moreover, our experimental results in Section 4 demonstrate that GD achieves linear convergence to critical points, which further supports our theoretical findings.
Third, we emphasize our technical contributions to establishing the error bound. Unlike prior works [30, 21, 22, 43], where the critical point set $\mathcal{W}_{F}$ has a relatively simple structure that allows for closed-form computation of the distance $\mathrm{dist}(\bm{W},\mathcal{W}_{F})$, the presence of orthogonal permutations and hierarchical structures in $\mathcal{W}_{F}$ (see 1) prevents such computation. To address this challenge, we develop a suite of new techniques (see Section 3.2.2), which offer a new framework for establishing error bounds of non-convex problems with complex solution sets.
3 Proofs of the Main Results
In this section, we present the proofs of our main results on the critical point set (i.e., 1) in Section 3.1 and the error bound (i.e., 2) for Problem (1) in Section 3.2. To facilitate readability, we summarize the notations used in this section in Table 1.
3.1 Proofs of the Critical Point Set
To characterize the critical points and establish the error bound for Problem (1), we claim that it suffices to study the following problem:
$\min_{\bm{W}}\ G(\bm{W}):=\left\|\bm{W}_{L}\cdots\bm{W}_{1}-\sqrt{\lambda}\bm{Y}\right\|_{F}^{2}+\lambda\sum_{l=1}^{L}\|\bm{W}_{l}\|_{F}^{2},$ (12)
where $\lambda:=\prod_{l=1}^{L}\lambda_{l}$, as defined in (7) and (8). We compute the gradient of $G(\bm{W})$ as follows:
$\nabla_{l}G(\bm{W})=2\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{l-1:1}^{T}+2\lambda\bm{W}_{l},\ \forall l\in[L].$ (13)
The critical point set of Problem (12) is defined as
$\mathcal{W}_{G}:=\left\{\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L}):\nabla G(\bm{W})=\bm{0}\right\}.$
Notation Eq. Notation Eq. Notation Eq. Notation Eq. Notation Eq. Notation Eq.
$p$ (45) $p_{Y},r_{Y}$ (4) $s_{i},h_{i}$ (4) $t_{i},g_{i}$ (45) $\varphi(x)$ (85) $G(\bm{W})$ (12)
$\lambda$ (7,8) $\delta_{y}$ (18) $\delta_{\sigma},\mathcal{\bm{Y}}$ (28) $\mathcal{\bm{A}}$ (26) $\mathcal{\bm{B}}$ (27) $\mathcal{\bm{A}}_{\text{sort}}$ (29)
$\eta_{1}$ (56) $\eta_{2}$ (84) $\eta_{3},\eta_{4}$ (88) $\eta_{5}$ (90) $L_{G}$ (144) $c_{1}$ (42)
$c_{2}$ (57) $c_{3}$ (74) $c_{4}$ (90) $c_{5}$ (95) $\delta_{1},\delta_{2}$ (87) $\sigma^{*}_{\min},\sigma_{\max}^{*},r_{\sigma}$ (34)
$\epsilon_{\bm{0}},\kappa_{\bm{0}}$ (31) $\epsilon_{\bm{\sigma}^{*}},\kappa_{\bm{\sigma}^{*}}$ (114) $\mathcal{W}_{\bm{\sigma}^{*}}$ (19,29) $\kappa,\epsilon$ (125) $\kappa_{1},\epsilon_{1}$ (126) $d_{\max}$ (18)
Table 1: The explicit expressions of the notations in the paper
Now, we present the following lemma to prove our claim.
Lemma 1.
Consider Problems (1) and (12). The following statements hold:
(i) $(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}_{F}$ if and only if $(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L})\in\mathcal{W}_{G}$.
(ii) Let $\bm{Y}=\bm{U}_{Y}\bm{\Sigma}_{Y}\bm{V}_{Y}^{T}$ be an SVD of $\bm{Y}$. Then $(\bm{W}_{1},\ldots,\bm{W}_{L})$ is a critical point of Problem (12) if and only if $(\bm{W}_{1}\bm{V}_{Y},\,\bm{W}_{2},\,\ldots,\,\bm{W}_{L-1},\,\bm{U}_{Y}^{T}\bm{W}_{L})$ is a critical point of $\min_{\bm{W}}\,\big\|\,\bm{W}_{L:1}-\sqrt{\lambda}\bm{\Sigma}_{Y}\big\|_{F}^{2}+\lambda\sum_{l=1}^{L}\|\bm{W}_{l}\|_{F}^{2}.$
(iii) Suppose that there exist constants $\epsilon,\kappa>0$ such that for all $\bm{Z}$ satisfying $\mathrm{dist}(\bm{Z},\mathcal{W}_{G})\leq\epsilon$, we have
$\mathrm{dist}(\bm{Z},\mathcal{W}_{G})\leq\kappa\|\nabla G(\bm{Z})\|_{F}.$ (14)
Then for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\epsilon/\sqrt{\lambda_{\max}}$, it holds that
$\mathrm{dist}\left(\bm{W},\mathcal{W}_{F}\right)\leq({\kappa\lambda}/{\lambda_{\min}})\|\nabla F(\bm{W})\|_{F}.$
Proof.
(i) and (ii) have been proved in [6, Lemma 3.1]. Now, we prove (iii). Using (i), we have
$\mathcal{W}_{G}=\left\{(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L}):(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}_{F}\right\}.$
Let $\bm{W}=\left(\bm{W}_{1},\dots,\bm{W}_{L}\right)$ be arbitrary and $\hat{\bm{W}}=(\sqrt{\lambda_{1}}\bm{W}_{1},\dots,\sqrt{\lambda_{L}}\bm{W}_{L})$. Let $(\sqrt{\lambda_{1}}\hat{\bm{W}}_{1}^{*},\dots,\sqrt{\lambda_{L}}\hat{\bm{W}}_{L}^{*})$ with $(\hat{\bm{W}}_{1}^{*},\dots,\hat{\bm{W}}_{L}^{*})\in\mathcal{W}_{F}$ be such that $\mathrm{dist}^{2}(\hat{\bm{W}},\mathcal{W}_{G})=\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}-\hat{\bm{W}}^{*}_{l}\|_{F}^{2}$, and let $(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{F}$ be such that $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F})=\sum_{l=1}^{L}\|\bm{W}_{l}-\bm{W}_{l}^{*}\|_{F}^{2}$. We have
$\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F})$ $=\sum_{l=1}^{L}\|\bm{W}_{l}-\bm{W}_{l}^{*}\|_{F}^{2}\geq\frac{\sum_{l=1}^{L}\|\sqrt{\lambda_{l}}\bm{W}_{l}-\sqrt{\lambda_{l}}\bm{W}_{l}^{*}\|_{F}^{2}}{\lambda_{\max}}$
$\geq\frac{1}{\lambda_{\max}}\mathrm{dist}^{2}(\hat{\bm{W}},\mathcal{W}_{G}).$ (15)
Using this and $\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\epsilon/\sqrt{\lambda_{\max}}$ yields $\mathrm{dist}(\hat{\bm{W}},\mathcal{W}_{G})\leq\epsilon$, and thus (14) holds. Next, we have
$\textstyle\mathrm{dist}^{2}(\hat{\bm{W}},\mathcal{W}_{G})\!=\!\!\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}-\hat{\bm{W}}^{*}_{l}\|_{F}^{2}\geq\!\lambda_{\min}\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{F}),$ (16)
[…]
$\qquad\qquad\qquad\qquad\ =\frac{\lambda}{\sqrt{\lambda_{l}}}\nabla_{l}F(\bm{W}),\ \forall l\in[L].$ (17)
Therefore, we have
[…]
Using (i) and (iii) of the above lemma, it suffices to establish an error bound for Problem (12) in the rest of this section. By Lemma 1(ii), we assume without loss of generality that $\bm{Y}=\bm{\Sigma}_{Y}$ from now on, where $\bm{\Sigma}_{Y}$ is defined in (3) and its diagonal entries are partitioned as in (4). Moreover, we define
$\delta_{y}:=\min\left\{|y_{s_{i}}-y_{s_{i+1}}|:i\in[p_{Y}]\right\},\quad d_{\max}:=\max\{d_{0},d_{1},\ldots,d_{L}\}.$ (18)
Using [6, Proposition 3.3] and Assumption 1, it is straightfoward to verify that the critical point set of Problem (12) admits the following closed-form expression:
Fact 1.
Suppose that Assumption 1 holds. The critical point set of Problem (12) is
$\mathcal{W}_{G}=\bigcup_{(\bm{\sigma},\bm{\Pi})\in\mathcal{B}}\mathcal{W}_{(\bm{\sigma},\bm{\Pi})},$ (19)
where
$\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}:=\left\{\bm{W}:\begin{array}[]{l}\bm{\Sigma}_{l}=\mathrm{BlkD}\left(\mathrm{diag}(\bm{\sigma}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\ \forall l\in[L],\\ \bm{W}_{1}=\bm{Q}_{2}\bm{\Sigma}_{1}\mathrm{BlkD}\left(\bm{\Pi},\bm{I}\right)\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ \bm{W}_{l}=\bm{Q}_{l+1}\bm{\Sigma}_{l}\bm{Q}_{l}^{T},\ l=2,\dots,L-1,\ \bm{Q}_{l}\in\mathcal{O}^{d_{l-1}},\ l=2,\dots,L,\\ \bm{W}_{L}=\mathrm{BlkD}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\mathrm{BlkD}\left(\bm{\Pi}^{T},\bm{I}\right)\bm{\Sigma}_{L}\bm{Q}_{L}^{T},\\ \bm{O}_{i}\in\mathcal{O}^{h_{i}},\ \forall i\in[p_{Y}],\ \bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\ \widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}}\end{array}\right\},$ (25)
$\mathcal{A}:=\left\{\bm{a}\in\mathbb{R}^{d_{\min}}:\ a_{i}^{2L-1}-\sqrt{\lambda}y_{i}a_{i}^{L-1}+\lambda a_{i}=0,\ a_{i}\geq 0,\ \forall i\in[d_{\min}]\right\},$ (26)
$\mathcal{B}:=\left\{(\bm{\sigma},\bm{\Pi})\in\mathbb{R}^{d_{\min}}\times\mathcal{P}^{d_{\min}}:\bm{a}\in\mathcal{A},\ \bm{\sigma}=\bm{\Pi}\bm{a},\ \sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{d_{\min}}\right\}.$ (27)
Note that the permutation matrix $\bm{\Pi}$ can be absorbed into $\bm{\Sigma}_{l}$ and $\bm{Q}_{l}$ to simplify the above expression. With this observation, and by invoking Lemma 1, we are now ready to prove 1 as follows:
Proof.
By (ii) of Lemma 1, without loss of generality, we may assume that $\bm{Y}=\bm{\Sigma}_{Y}$, i.e., $\bm{U}_{Y}=\bm{I}$ and $\bm{V}_{Y}=\bm{I}$. Let $(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}_{F}$ be arbitrary. It follows from 1 and Lemma 1 that there exists $(\bm{\sigma},\bm{\Pi})\in\cal$ B such that $(\bm{W}_{1},\dots,\bm{W}_{L})$ takes a form of (25) with $\bm{\Sigma}_{l}=\mathrm{BlkDiag}(\mathrm{diag}(\bm{\sigma})/\sqrt{\lambda_{l}},\bm{0})$. According to (26) and (27), there exists $\bm{a}\in\cal A$ such that $\bm{\Pi}\mathrm{diag}(\bm{a})\bm{\Pi}^{T}=\mathrm{diag}(\bm{\sigma})$. Then, let
$\bm{\Sigma}_{l}^{\prime}:=\mathrm{BlkDiag}(\bm{\Pi}^{T},\bm{I}_{d_{l}-d_{\min}})\bm{\Sigma}_{l}\mathrm{BlkDiag}(\bm{\Pi},\bm{I}_{d_{l-1}-d_{\min}})=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{a})/\sqrt{\lambda_{l}},\bm{0}\right),\ \forall l\in[L],$
$\bm{Q}_{l}^{\prime}:=\bm{Q}_{l}\mathrm{BlkDiag}(\bm{\Pi},\bm{I}_{d_{l-1}-d_{\min}})\in\mathcal{O}^{d_{l-1}},\ \forall l\in[L].$
Therefore, we obtain that $(\bm{W}_{1},\dots,\bm{W}_{L})$ can be rewriten in the form of (5) with $\bm{\Sigma}_{l}=\bm{\Sigma}_{l}^{\prime}$, $\bm{Q}_{l}=\bm{Q}_{l}^{\prime}$. Conversely, let $(\bm{W}_{1},\dots,\bm{W}_{L})$ be of the form of (5) for some $\bm{\sigma}=(\sigma_{1},\ldots,\sigma_{r_{Y}},0,\ldots 0)\in\mathbb{R}^{d_{\min}}$. Then, we choose a permutation matrix $\bm{\Pi}\in\mathcal{P}^{d_{\min}}$ such that $\bm{\sigma}^{\prime}=\bm{\Pi}\bm{\sigma}\ \text{satisfies}\ \sigma^{\prime}_{1}\geq\sigma^{\prime}_{2}\geq\dots\geq\sigma^{\prime}_{d_{\min}}.$ Let
$\bm{\Sigma}^{\prime}_{l}:=\mathrm{BlkDiag}(\bm{\Pi},\bm{I}_{d_{l}-d_{\min}})\bm{\Sigma}_{l}\mathrm{BlkDiag}(\bm{\Pi}^{T},\bm{I}_{d_{l-1}-d_{\min}})=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{\sigma}^{\prime})/\sqrt{\lambda_{l}},\bm{0}\right),\ \forall l\in[L],$
$\bm{Q}^{\prime}_{l}:=\bm{Q}_{l}\mathrm{BlkDiag}(\bm{\Pi}^{T},\bm{I}_{d_{l-1}-d_{\min}}),\ \forall l\in[L].$
Then, we obtain that $(\bm{W}_{1},\dots,\bm{W}_{L})$ can be rewriten in the form of (25) with $\bm{\Sigma}_{l}=\bm{\Sigma}_{l}^{\prime}$, $\bm{Q}_{l}=\bm{Q}_{l}^{\prime}$. This, together with 1 and item (i) and (ii) of Lemma 1, implies $(\bm{W}_{1},\dots,\bm{W}_{L})\in\mathcal{W}_{F}$. ∎
3.2 Analysis of the Error Bound
In this subsection, we establish the error bound for Problem (12) based on 1. In Section 3.2.1, we first present preliminary results for our analysis. In Section 3.2.2, we provide the main proofs of the error bound. For ease of exposition, we define
$\mathcal{Y}:=\bigcup_{i\in[d_{\min}]}\left\{\sigma\geq 0:\sigma^{2L-1}+\lambda\sigma-\sqrt{\lambda}y_{i}\sigma^{L-1}=0\right\},\ \delta_{\sigma}:=\min\left\{|x-y|:x\neq y\in\mathcal{Y}\right\}.$ (28)
The following result elucidates the structure of the collection $\{\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}\}_{(\bm{\sigma},\bm{\Pi})\in\mathcal{B}}$.
Proposition 1.
Let $(\bm{\sigma},\bm{\Pi})\in\mathcal{B}$ and $(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})\in\mathcal{B}$ be arbitrary. The following statements hold:
(i) It holds that $\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}=\mathcal{W}_{(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})}$ if and only if $\bm{\sigma}=\bm{\sigma}^{\prime}$.
(ii) If $\bm{\sigma}\neq\bm{\sigma}^{\prime}$, it holds that $\mathrm{dist}\left(\mathcal{W}_{(\bm{\sigma},\bm{\Pi})},\mathcal{W}_{(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})}\right)\geq\delta_{\sigma}.$
We defer the proof of this proposition to Appendix B.1. This proposition demonstrates that for any pair $(\bm{\sigma},\bm{\Pi})\in\mathcal{B}$ and $(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})\in\mathcal{B}$, if $\bm{\sigma}\neq\bm{\sigma}^{\prime}$, $\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}$ is well separated from $\mathcal{W}_{(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})}$, and otherwise they are identical. Therefore, with a slight abuse of notation, we write $\mathcal{W}_{\bm{\sigma}}:=\mathcal{W}_{(\bm{\sigma},\bm{\Pi})}$ for any $(\bm{\sigma},\bm{\Pi})\in\mathcal{B}$. Thus, we can express the set of critical points of Problem (12) as follows:
$\mathcal{W}_{G}:=\bigcup_{\bm{\sigma}\in\mathcal{A}_{\rm sort}}\mathcal{W}_{\bm{\sigma}},\ \text{where}\ \mathcal{A}_{\rm sort}:=\left\{\bm{\sigma}\in\mathbb{R}^{d_{\min}}:(\bm{\sigma},\bm{\Pi})\in\mathcal{B}\right\}.$ (29)
3.2.1 Preliminary Results
According to 1 and (29), for any $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{G})\leq\delta_{\sigma}/2$, there exists a $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}$ such that $\mathrm{dist}(\bm{W},\mathcal{W}_{G})=\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}}).$ This observation simplifies our analysis, as it suffices to bound $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})$ for each $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}$. Note that $\bm{0}\in\mathcal{A}_{\rm sort}$. For this case, we directly show the error bound as follows:
Proposition 2.
Suppose that Assumptions 1 and 2 hold. Then for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{0}})\leq\epsilon_{\bm{0}}$, we have
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{0}})\leq\kappa_{\bm{0}}\,\bigl\|\nabla G(\bm{W})\bigr\|_{F},$ (30)
where
$(\epsilon_{\bm{0}},\,\kappa_{\bm{0}}):=\begin{cases}\Biggl(\Bigl(\frac{\sqrt{\lambda}}{2(\sqrt{\lambda}+y_{1})}\min\!\Bigl\{\min\limits_{i\in[s_{p_{Y}}]}\bigl|\lambda-y_{i}^{2}\bigr|,\,\lambda\Bigr\}\Bigr)^{\frac{1}{2}},\;\frac{2(\sqrt{\lambda}+y_{1})}{\sqrt{\lambda}\min\!\Bigl\{\min\limits_{i\in[s_{p_{Y}}]}\bigl|\lambda-y_{i}^{2}\bigr|,\,\lambda\Bigr\}}\Biggr),\!\!&\!\!\!\!L=2,\\ \Biggl(\min\!\left\{\left(\frac{\lambda}{3}\right)^{{1}/{(2L-2)}},\;\left(\frac{\sqrt{\lambda}}{3y_{1}}\right)^{{1}/{(L-2)}}\right\},\;\frac{3\sqrt{L}}{2\lambda}\Biggr),\!\!&\!\!\!\!L\geq 3.\end{cases}$ (31)
Proof.
Let’s begin by focusing on $L=2$. Using 1, we have $\mathcal{W}_{\bm{0}}=\{(\bm{0},\bm{0})\}$ and $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{\bm{0}})=\|\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{2}\|_{F}^{2}.$ Without loss of generality, assume that $\|\bm{W}_{1}\|_{F}\geq\|\bm{W}_{2}\|_{F}$. We have
$\nabla_{1}G(\bm{W})=2\bm{W}_{2}^{T}\left(\bm{W}_{2}\bm{W}_{1}-\sqrt{\lambda}\bm{Y}\right)+2\lambda\bm{W}_{1},$
$\nabla_{2}G(\bm{W})=2\left(\bm{W}_{2}\bm{W}_{1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{1}^{T}+2\lambda\bm{W}_{2}.$ (32)
Using the first equation in (32), we have
$\frac{\sqrt{\lambda}}{2}\|\nabla_{1}G(\bm{W})\|_{F}$ $\geq\left\|\lambda\bm{W}_{2}^{T}\bm{Y}-\lambda^{\frac{3}{2}}\bm{W}_{1}\right\|_{F}-\sqrt{\lambda}\|\bm{W}_{2}^{T}\bm{W}_{2}\bm{W}_{1}\|_{F}$
$\geq\left\|\lambda\bm{W}_{2}^{T}\bm{Y}-\lambda^{\frac{3}{2}}\bm{W}_{1}\right\|_{F}-\sqrt{\lambda}\|\bm{W}_{2}\|_{F}^{2}\|\bm{W}_{1}\|_{F}.$ (33)
Multiplying $\bm{Y}^{T}$ on the both sides of the second equation in (32), together with the triangle inequality, yields
$\frac{1}{2}\left\|\bm{Y}^{T}\nabla_{2}G(\bm{W})\right\|_{F}\geq\|\sqrt{\lambda}\bm{Y}^{T}\bm{Y}\bm{W}_{1}^{T}-\lambda\bm{Y}^{T}\bm{W}_{2}\|_{F}-\|\bm{Y}^{T}\bm{W}_{2}\bm{W}_{1}\bm{W}_{1}^{T}\|_{F},$
which implies
$\frac{y_{1}}{2}\|\nabla_{2}G(\bm{W})\|_{F}\geq\|\sqrt{\lambda}\bm{Y}^{T}\bm{Y}\bm{W}_{1}^{T}-\lambda\bm{Y}^{T}\bm{W}_{2}\|_{F}-y_{1}\|\bm{W}_{2}\|_{F}\|\bm{W}_{1}\|_{F}^{2}.$
Summing up this inequality with (3.2.1) and using the triangle inequality, we obtain
$\frac{\sqrt{\lambda}+y_{1}}{2}\|\nabla G(\bm{W})\|_{F}$ $\geq\frac{\sqrt{\lambda}}{2}\|\nabla_{1}G(\bm{W})\|_{F}+\frac{y_{1}}{2}\left\|\nabla_{2}G(\bm{W})\right\|_{F}$
$\geq\sqrt{\lambda}\left\|\bm{W}_{1}\left(\lambda\bm{I}-\bm{Y}^{T}\bm{Y}\right)\right\|_{F}-\|\bm{W}_{1}\|_{F}\|\bm{W}_{2}\|_{F}\left(y_{1}\|\bm{W}_{1}\|_{F}+\sqrt{\lambda}\|\bm{W}_{2}\|_{F}\right)$
$\geq\sqrt{\lambda}\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\lambda-y_{i}^{2}\right|,\lambda\right\}\|\bm{W}_{1}\|_{F}-\left(\sqrt{\lambda}+y_{1}\right)\|\bm{W}_{1}\|_{F}^{3}$
${\geq}\frac{\sqrt{\lambda}}{2}\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\lambda-y_{i}^{2}\right|,\lambda\right\}\|\bm{W}_{1}\|_{F},$
$\geq\frac{\sqrt{\lambda}}{4}\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\lambda-y_{i}^{2}\right|,\lambda\right\}\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{0}})$
where the third inequality follows from (7) and $\bm{Y}=\bm{\Sigma}_{Y}$ defined in (3), $\|\bm{A}\bm{B}\|_{F}\geq\sigma_{\min}(\bm{A})\|\bm{B}\|_{F}$, and $\|\bm{W}_{1}\|_{F}\geq\|\bm{W}_{2}\|_{F}$, and the fourth inequality uses $\|\bm{W}\|_{F}\leq\epsilon_{\bm{0}}$ and (31). This directly implies (30).
Next, we focus on $L\geq 3$. According to 1 and (29), we have $\mathcal{W}_{\bm{0}}=\{(\bm{0},\bm{0},\dots,\bm{0})\}$. Then, we have $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{\bm{0}})=\sum_{l=1}^{L}\|\bm{W}_{l}\|_{F}^{2}\leq\epsilon_{\bm{0}}.$ Let $k\in[L]$ be such that $k\in\operatorname{\arg\max}_{l\in[L]}\|\bm{W}_{l}\|_{F}$. Then, we compute
[…]
$\geq\lambda\|\bm{W}_{k}\|_{F}-\left(\|\bm{W}_{k}\|_{F}^{2L-1}+\sqrt{\lambda}y_{1}\|\bm{W}_{k}\|_{F}^{L-1}\right)$
[…]
Thus, we have $\mathrm{dist}^{2}(\bm{W},\mathcal{W}_{\bm{0}})\leq L\|\bm{W}_{k}\|_{F}^{2}\leq\frac{9L}{4\lambda^{2}}\left\|\nabla_{k}G(\bm{W})\right\|_{F}^{2},$ which directly implies the desired result. ∎
According to this proposition, it suffices to consider $\bm{\sigma}^{*}\neq\bm{0}$ from now on. Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary and define
$\sigma_{\min}^{*}:=\min\left\{\sigma_{i}^{*}\neq 0:i\in[d_{\min}]\right\},\ \sigma_{\max}^{*}:=\max\left\{\sigma_{i}^{*}:i\in[d_{\min}]\right\},\ \text{and}\ \ r_{\sigma}:=\|\bm{\sigma}^{*}\|_{0}.$ (34)
By the definition of $\delta_{\sigma}$ in (28) and $\bm{0}\in\mathcal{Y}$, we have $\delta_{\sigma}\leq\sigma^{*}_{\min}$ for each $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}$. Then, we show the following results.
Lemma 2.
Suppose that Assumption 1 holds. Let $\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L})$ and $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})<{\sigma^{*}_{\min}}/{2}$. Then the following statements hold:
$\text{(i)}\quad\frac{\sigma_{\max}^{*}}{2}\leq\|\bm{W}_{l}\|\leq\frac{3\sigma_{\max}^{*}}{2},\ \sigma_{i}(\bm{W}_{l})\geq\frac{\sigma_{\min}^{*}}{2},\ \forall i\in[r_{\sigma}],\ l\in[L].$ (35)
$\text{(ii)}\quad\|\bm{W}_{l+1}^{T}\bm{W}_{l+1}-\bm{W}_{l}\bm{W}_{l}^{T}\|_{F}\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda}\|\nabla G(\bm{W})\|_{F},\ \forall l\in[L-1].$ (36)
$\text{(iii)}\quad|\sigma_{i}(\bm{W}_{l})-\sigma_{i}(\bm{W}_{l+1})|\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F},\ \forall i\in[r_{\sigma}],\ l\in[L-1].$ (37)
$\text{(iv)}\quad\left\|\bm{W}_{j:i}^{T}\bm{W}_{j:i}-(\bm{W}^{T}_{i}\bm{W}_{i})^{j-i+1}\right\|_{F}\leq\frac{(j-i)(j-i+1)}{2\sqrt{2}\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2j-2i+1}\|\nabla G(\bm{W})\|_{F},\ \forall i\leq j,$ (38)
$\qquad\ \ \left\|\bm{W}_{i:j}\bm{W}_{i:j}^{T}-(\bm{W}_{i}\bm{W}_{i}^{T})^{i-j+1}\right\|_{F}\leq\frac{(i-j)(i-j+1)}{2\sqrt{2}\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2i-2j+1}\|\nabla G(\bm{W})\|_{F},\ \forall i>j.$ (39)
Proof.
(i) Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. Using the triangle inequality, we have for each $l\in[L]$,
$\|\bm{W}_{l}\|\geq\|\bm{W}_{l}^{*}\|-\|\bm{W}_{l}-\bm{W}_{l}^{*}\|\geq\|\bm{W}_{l}^{*}\|-\|\bm{W}_{l}-\bm{W}_{l}^{*}\|_{F}\geq\sigma^{*}_{\max}-\frac{\sigma^{*}_{\min}}{2}\geq\frac{\sigma^{*}_{\max}}{2},$
where the third inequality follows from 1 and $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})<{\sigma^{*}_{\min}}/{2}$. Similarly, we have $\|\bm{W}_{l}\|\leq 3\sigma^{*}_{\max}/2$ for each $l\in[L]$. Using Weyl’s inequality (see, e.g., [18, Corollary 7.3.5]), we have $\left|\sigma_{i}(\bm{W}_{l})-\sigma_{i}^{*}\right|\leq\|\bm{W}_{l}-\bm{W}_{l}^{*}\|\leq\sigma_{\min}^{*}/{2}$ for all $i\in[r_{\sigma}]$ and $l\in[L]$. This implies $\sigma_{i}(\bm{W}_{l})\geq\sigma_{i}^{*}-{\sigma_{\min}^{*}}/{2}\geq{\sigma_{\min}^{*}}/{2}.$
(ii) According to (13), we compute
$\nabla_{l}G(\bm{W})\bm{W}_{l}^{T}=2\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{l:1}^{T}+2\lambda\bm{W}_{l}\bm{W}_{l}^{T},$
$\bm{W}_{l+1}^{T}\nabla_{l+1}G(\bm{W})=2\bm{W}_{L:l+1}^{T}\left(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y}\right)\bm{W}_{l:1}^{T}+2\lambda\bm{W}_{l+1}^{T}\bm{W}_{l+1}.$
Then, we have $\left(\nabla_{l}G(\bm{W})\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\nabla_{l+1}G(\bm{W})\right)/2=\lambda\left(\bm{W}_{l}\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\bm{W}_{l+1}\right)$ for each $l\in[L-1]$. This implies
$\left\|\bm{W}_{l}\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\bm{W}_{l+1}\right\|_{F}$ $=\frac{1}{2\lambda}\left\|\nabla_{l}G(\bm{W})\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\nabla_{l+1}G(\bm{W})\right\|_{F}$
$\leq\frac{3\sigma_{\max}^{*}}{4\lambda}\left(\|\nabla_{l}G(\bm{W})\|_{F}+\|\nabla_{l+1}G(\bm{W})\|_{F}\right)\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda}\|\nabla G(\bm{W})\|_{F},$
where the first inequality follows from the triangle inequality, $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|\|\bm{B}\|_{F}$, and (35).
(iii) Using Weyl’s inequality and (36), we obtain that for each $l\in[L]$ and $i\in[r_{\sigma}]$,
$\left|\sigma_{i}^{2}(\bm{W}_{l})-\sigma_{i}^{2}(\bm{W}_{l+1})\right|\leq\left\|\bm{W}_{l}\bm{W}_{l}^{T}-\bm{W}_{l+1}^{T}\bm{W}_{l+1}\right\|_{F}\leq\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda}\|\nabla G(\bm{W})\|_{F}.$
This implies
[…]
(iv) We first prove (38). Note that
$\textstyle\bm{W}_{j:i}^{T}\bm{W}_{j:i}-(\bm{W}^{T}_{i}\bm{W}_{i})^{j-i+1}=\sum_{k=1}^{j-i}\bm{W}_{j-k:i}^{T}\left((\bm{W}_{j-k+1}^{T}\bm{W}_{j-k+1})^{k}-(\bm{W}_{j-k}\bm{W}_{j-k}^{T})^{k}\right)\bm{W}_{j-k:i}.$ (40)
For each $k\in[j-i]$, we compute
$\left\|\bm{W}_{j-k:i}^{T}\left((\bm{W}_{j+1-k}^{T}\bm{W}_{j+1-k})^{k}-(\bm{W}_{j-k}\bm{W}_{j-k}^{T})^{k}\right)\bm{W}_{j-k:i}\right\|_{F}$
$\leq$ $\left\|(\bm{W}_{j+1-k}^{T}\bm{W}_{j+1-k})^{k}-(\bm{W}_{j-k}\bm{W}_{j-k}^{T})^{k}\right\|_{F}\prod_{l=i}^{j-k}\|\bm{W}_{l}\|^{2}$
$\leq$ $k\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(j-i)}\left\|\bm{W}_{j+1-k}^{T}\bm{W}_{j+1-k}-\bm{W}_{j-k}\bm{W}_{j-k}^{T}\right\|_{F}$
[…]
$=\frac{\sqrt{2}k}{2\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(j-i)+1}\|\nabla G(\bm{W})\|_{F},$
where the second inequality follows from
$\left\|(\bm{W}_{j+1-k}^{T}\bm{W}_{j+1-k})^{k}-(\bm{W}_{j-k}\bm{W}_{j-k}^{T})^{k}\right\|_{F}$
[…]
This, together with (40), yields
$\left\|\bm{W}_{j:i}^{T}\bm{W}_{j:i}-(\bm{W}^{T}_{i}\bm{W}_{i})^{j-i+1}\right\|_{F}$ $\leq\left(\sum_{k=1}^{j-i}k\right)\frac{\sqrt{2}}{2\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2j-2i+1}\|\nabla G(\bm{W})\|_{F}$
$\leq\frac{(j-i)(j-i+1)}{2\sqrt{2}\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2j-2i+1}\|\nabla G(\bm{W})\|_{F}.$
Using the same argument, we can prove (39). ∎
Corollary 2.
Under the setting of Lemma 2, we have
$\left\|(\bm{W}_{l}\bm{W}^{T}_{l})^{L-1}\bm{W}_{l}-\sqrt{\lambda}\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}^{T}+\lambda\bm{W}_{l}\right\|_{F}\leq c_{1}\|\nabla G(\bm{W})\|_{F},\ \forall l\in[L],$ (41)
where
$c_{1}:=\max_{l\in[L]}\left\{\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}\frac{(L-l)(L-l+1)+(l-1)l}{2\sqrt{2}\lambda}+\frac{1}{2}\right\}.$ (42)
Proof.
For ease of exposition, let
$\bm{R}_{1}(l):=\bm{W}_{l-1:1}\bm{W}_{l-1:1}^{T}-(\bm{W}_{l}^{T}\bm{W}_{l})^{l-1},\ l=2,3,\dots,L,$ $\bm{R}_{1}(1):=\bm{0},$
$\bm{R}_{2}(l):=\bm{W}_{L:l+1}^{T}\bm{W}_{L:l+1}-(\bm{W}_{l}\bm{W}_{l}^{T})^{L-l},\ l=1,2,\dots,L-1,$ $\bm{R}_{2}(L):=\bm{0}.$
For each $l\in[L]$, we compute
$\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T})^{l-1}-(\bm{W}_{l}^{T}\bm{W}_{l})^{l-1}\right\|_{F}$
$\leq\quad$ $\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T})^{l-2}(\bm{W}_{l-1}\bm{W}_{l-1}^{T}-\bm{W}_{l}^{T}\bm{W}_{l})\right\|_{F}+\cdots+\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T}-\bm{W}_{l}^{T}\bm{W}_{l})(\bm{W}_{l}^{T}\bm{W}_{l})^{l-2}\right\|_{F}$
[…]
This, together with the triangle inequality and (39), yields that for $l=2,3,\dots,L$,
$\|\bm{R}_{1}(l)\|_{F}\leq$ $\ \left\|\bm{W}_{l-1:1}\bm{W}_{l-1:1}^{T}\!-\!(\bm{W}_{l-1}\bm{W}_{l-1}^{T})^{l-1}\right\|_{F}\!+\!\left\|(\bm{W}_{l-1}\bm{W}_{l-1}^{T})^{l-1}\!\!\!-\!(\bm{W}_{l}^{T}\bm{W}_{l})^{l-1}\right\|_{F}$
$\leq$ $\ \frac{l(l-1)}{2\sqrt{2}\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2l-3}\|\nabla G(\bm{W})\|_{F}.$ (43)
Using the same argument and (38), we have for $l=1,2,\dots,L-1$,
$\|\bm{R}_{2}(l)\|_{F}\leq\frac{(L-l+1)(L-l)}{2\sqrt{2}\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(L-l)-1}\|\nabla G(\bm{W})\|_{F}.$ (44)
For each $l=1,2,\ldots L$, we compute
[…]
$=$~ $(\bm{W}_{l}\bm{W}^{T}_{l})^{L-1}\bm{W}_{l}-\sqrt{\lambda}\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}^{T}$
$+\lambda\bm{W}_{l}+\bm{R}_{2}(l)\bm{W}_{l}\bm{W}_{l-1:1}\bm{W}_{l-1:1}^{T}+\left(\bm{W}_{l}\bm{W}_{l}^{T}\right)^{L-l}\bm{W}_{l}\bm{R}_{1}(l).$
This, together with the triangle inequality, (35), (3.2.1), and (44), yields for each $l\in[L]$,
$\left\|(\bm{W}_{l}\bm{W}^{T}_{l})^{L-1}\bm{W}_{l}-\sqrt{\lambda}\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}^{T}+\lambda\bm{W}_{l}\right\|_{F}$
[…]
$\leq\$$\left(\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}\frac{(L-l)(L-l+1)+(l-1)l}{2\sqrt{2}\lambda}+\frac{1}{2}\right)\|\nabla G(\bm{W})\|_{F},$
which directly implies (41). ∎
To handle the repeated singular values in $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$, let $p\geq 1$ denote the number of distinct positive singular values. Recall from (34) that $r_{\sigma}$ denotes the number of positive singular values of $\bm{\sigma}^{*}$. Then there exist indices $t_{0},t_{1},\cdots,t_{p}$ such that $0=t_{0}<t_{1}<\cdots<t_{p}=r_{\sigma}$ and
$\sigma_{t_{0}+1}^{*}=\dots=\sigma_{t_{1}}^{*}>\sigma_{t_{1}+1}^{*}=\dots=\sigma_{t_{2}}^{*}>\dots>\sigma_{t_{p-1}+1}^{*}=\dots=\sigma_{t_{p}}^{*}>0.$ (45)
Let $g_{i}:=t_{i}-t_{i-1}$ be the multiplicity of the $i$-th largest positive value of $\bm{\sigma}^{*}$ for each $i\in[p]$ and $g_{\max}:=\max\{g_{i}:i\in[p]\}$. Obviously, we have $r_{\sigma}=\sum_{i=1}^{p}g_{i}$ . Moreover, let
$\bm{W}_{l}$ $=\bm{U}_{l}\bm{\Sigma}_{l}\bm{V}_{l}^{T}$
$=\begin{bmatrix}\bm{U}_{l}^{(1)}&\!\!\!\!\dots\!\!\!&\bm{U}_{l}^{(p)}&\!\!\!\!\bm{U}_{l}^{(p+1)}\end{bmatrix}\mathrm{BlkD}\left(\bm{\Sigma}_{l}^{(1)},\dots,\bm{\Sigma}_{l}^{(p)},\bm{\Sigma}_{l}^{(p+1)}\right){\begin{bmatrix}\bm{V}_{l}^{(1)}&\!\!\!\!\!\!\dots\!\!\!\!\!&\!\bm{V}_{l}^{(p)}&\!\!\!\!\!\!\bm{V}_{l}^{(p+1)}\end{bmatrix}}^{T}$ (46)
be an SVD of $\bm{W}_{l}$ for each $l\in[L]$, where $\bm{U}_{l}\in\mathcal{O}^{d_{l}}$ with $\bm{U}_{l}^{(i)}\in\mathbb{R}^{d_{l}\times g_{i}}$ for each $i\in[p]$ and $\bm{U}_{l}^{(p+1)}\in\mathbb{R}^{d_{l}\times(d_{l}-r_{\sigma})}$, $\bm{\Sigma}_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ with decreasing singular values, $\bm{\Sigma}_{l}^{(i)}\in\mathbb{R}^{g_{i}\times g_{i}}$ for each $i\in[p]$, and $\bm{\Sigma}_{l}^{(p+1)}\in\mathbb{R}^{(d_{l}-r_{\sigma})\times(d_{l-1}-r_{\sigma})}$, and $\bm{V}_{l}\in\mathcal{O}^{d_{l-1}}$ with $\bm{V}_{l}^{(i)}\in\mathbb{R}^{d_{l-1}\times g_{i}}$ for each $i\in[p]$ and $\bm{V}_{l}^{(p+1)}\in\mathbb{R}^{d_{l-1}\times(d_{l-1}-r_{\sigma})}$.
With the above setup, we are ready to show that, for any point in the neighborhood of the set of critical points, the product of $\bm{U}_{l-1}$ and $\bm{V}_{l}$ is close to a block diagonal matrix with orthogonal diagonal blocks. Recall that $\delta_{\sigma}$ and $\mathcal{A}_{\rm sort}$ are defined in (28) and (29), respectively.
Proposition 3.
Suppose that Assumption 1 holds. Let $\bm{W}=(\bm{W}_{1},\dots,\bm{W}_{L})$ and $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary such that
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\frac{\delta_{\sigma}}{3}.$ (47)
For each $l\in\{2,\dots,L\}$, there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for all $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that
$\left\|\bm{U}_{l-1}^{T}\bm{V}_{l}-\mathrm{BlkD}\left(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}\right)\right\|_{F}\leq\frac{9\sigma_{\max}^{*}}{4\delta_{\sigma}\lambda\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F}.$ (48)
where $\bm{U}_{l}$ and $\bm{V}_{l}$ for each $l\in[L]$ are defined in (3.2.1).
Proof.
Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. For ease of exposition, let $\bm{H}_{l}:=\bm{U}_{l-1}^{T}\bm{V}_{l}$ for each $l\in\{2,\dots,L\}$. According to (3.2.1), the $(i,j)$-th block of $\bm{H}_{l}$, denoted by $\bm{H}_{l}^{(i,j)}$, is
$\bm{H}_{l}^{(i,j)}:=\bm{U}_{l-1}^{(i)^{T}}\bm{V}_{l}^{(j)},\ \forall(i,j)\in[p+1]\times[p+1].$ (49)
By (47) with $\delta_{\sigma}\leq\sigma_{\min}^{*}$ and Lemma 2, we have (36). Substituting (3.2.1) and $\bm{H}_{l}=\bm{U}_{l-1}^{T}\bm{V}_{l}$ into (36) yields
$\left\|\bm{H}_{l}\bm{\Sigma}_{l}^{T}\bm{\Sigma}_{l}-\bm{\Sigma}_{l-1}\bm{\Sigma}_{l-1}^{T}\bm{H}_{l}\right\|_{F}\leq\frac{3\sqrt{2}\sigma^{*}_{\max}}{4\lambda}\|\nabla G(\bm{W})\|_{F},\ \forall l=2,\dots,L.$
Using this and the block diagonal structures of $\bm{\Sigma}_{l}$ in (3.2.1) and $\bm{H}_{l}$ in (49), we obtain
$\ \sum_{j=1}^{p+1}\sum_{i=1,i\neq j}^{p+1}\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}-\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}^{2}$
$\leq$ $\ \sum_{j=1}^{p+1}\sum_{i=1}^{p+1}\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}-\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}^{2}\leq\left(\frac{3\sqrt{2}\sigma^{*}_{\max}}{4\lambda}\right)^{2}\|\nabla G(\bm{W})\|_{F}^{2}.$ (50)
[…]
$\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}\right\|_{F}\geq\lambda_{\min}^{2}(\bm{\Sigma}_{l}^{(j)})\|\bm{H}_{l}^{(i,j)}\|_{F}\geq\left(\sigma^{*}_{t_{j}}-\frac{\delta_{\sigma}}{3}\right)^{2}\|\bm{H}_{l}^{(i,j)}\|_{F},$ (51)
$\left\|\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}^{2}\leq\lambda_{\max}\left(\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\right)\|\bm{H}_{l}^{(i,j)}\|_{F}\leq\left(\sigma^{*}_{t_{i}}+\frac{\delta_{\sigma}}{3}\right)^{2}\|\bm{H}_{l}^{(i,j)}\|_{F}.$ (52)
For each $l\in[L]$ and $i>j\in[p+1]$, we bound
$\quad\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}-\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}\geq\left\|\bm{H}_{l}^{(i,j)}\bm{\Sigma}_{l}^{(j)^{T}}\bm{\Sigma}_{l}^{(j)}\right\|_{F}-\left\|\bm{\Sigma}_{l-1}^{(i)}\bm{\Sigma}_{l-1}^{(i)^{T}}\bm{H}_{l}^{(i,j)}\right\|_{F}$
[…]
Applying the same argument to the case where $i<j$, we obtain the same result as above. This, togher with (3.2.1), yields the following bound for off-diagonal blocks of $\bm{H}_{l}$:
$\sum_{j=1}^{p+1}\sum_{i=1,i\neq j}^{p+1}\|\bm{H}_{l}^{(i,j)}\|_{F}^{2}\leq\frac{81\sigma_{\max}^{*2}}{32\delta_{\sigma}^{2}\lambda^{2}\sigma^{*2}_{\min}}\|\nabla G(\bm{W})\|_{F}^{2}.$ (53)
For each $l\in[L]$ and $i\in[p+1]$, let $\bm{H}_{l}^{(i,i)}=\bm{P}_{l}^{(i)}\bm{\Lambda}_{l}^{(i)}\bm{Q}_{l}^{(i)^{T}}$ be a full SVD of $\bm{H}_{l}^{(i,i)}$, where $\bm{P}_{l}^{(i)},\ \bm{Q}_{l}^{(i)}$ are square orthogonal matrix. Then, we compute
$\left\|\bm{H}_{l}^{(i,i)}-\bm{P}_{l}^{(i)}\bm{Q}_{l}^{(i)^{T}}\right\|_{F}^{2}$ $=\sum_{k=1}^{g_{i}}\left(1-\sigma_{k}(\bm{H}_{l}^{(i,i)})\right)^{2}\leq\sum_{k=1}^{g_{i}}\left(1-\sigma_{k}^{2}(\bm{H}_{l}^{(i,i)})\right)^{2}$
$=\left\|\bm{I}-\bm{H}_{l}^{(i,i)}\bm{H}_{l}^{(i,i)^{T}}\right\|_{F}^{2}=\sum_{j=1,j\neq i}^{p+1}\|\bm{H}_{l}^{(i,j)}\bm{H}_{l}^{(i,j)^{T}}\|_{F}^{2}\leq\sum_{j=1,j\neq i}^{p+1}\|\bm{H}_{l}^{(i,j)}\|_{F}^{2},$
where the last equality is due to $\bm{H}_{l}^{(i,1)},\dots,\bm{H}_{l}^{(i,p+1)}]=\bm{U}_{l-1}^{(i)^{T}}\bm{V}_{l}$ with $\bm{U}_{l-1},\ \bm{V}_{l}\in\mathcal{O}^{d_{l-1}}$, and last inequality follows from $\|\bm{H}_{l}^{(i,j)}\|\leq 1$ for all $l\in[L]$ and $i,j\in[p+1]$. Therefore, we obtain
[…]
This, together with the definition of $\bm{H}_{l}$, (53) and $\bm{T}_{l}^{(i)}=\bm{P}_{l}^{(i)}\bm{Q}_{l}^{(i)^{T}}$, yields (48). ∎
Next, using the above proposition, we show that for weight matrices in the neighborhood of the set of critical points, the singular matrices $\bm{\Sigma}_{1}$ and $\bm{\Sigma}_{L}$ satisfy the following spectral inequalities.
Lemma 3.
Consider the setting of Lemma 2. Suppose that there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for all $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (48) holds for all $l=2,\dots,L$. Let $c_{1}$ be defined in (42) and
$\textstyle\bm{A}_{i}:=\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{1}^{(i)}\right)^{L-1},\ \forall i\in[p],\ \bm{A}_{p+1}:=\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(p+1)}\bm{\Sigma}_{l+1}^{(p+1)^{T}},$ (54)
$\textstyle\bm{B}_{i}:=\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1},\ \forall i\in[p],\ \bm{B}_{p+1}:=\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(p+1)^{T}}\bm{T}_{l+1}^{(p+1)},$ (55)
$\eta_{1}:=\frac{\sigma_{\max}^{*}}{\sigma_{\min}^{*}}\left(\frac{3\sqrt{2}\sigma_{\min}^{*}}{4\lambda}+\frac{81\sigma_{\max}^{*{2}}}{8\delta_{\sigma}\lambda}+\frac{9\sqrt{2g_{\max}}L\sigma_{\max}^{*}}{4\lambda}\right),$ (56)
$c_{2}:=\left(\frac{3}{2}\sigma_{\max}^{*}\right)^{L}\frac{3y_{1}L}{2\sqrt{\lambda}\delta_{\sigma}\sigma^{*}_{\min}}+c_{1}+y_{1}p\sqrt{\lambda}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{L-2}$
$\qquad\ \left(\frac{L^{2}\eta_{1}}{2\sigma^{*}_{\min}}+\frac{3\sqrt{2g_{\max}}L^{2}\sigma_{\max}^{*}}{2\lambda\sigma_{\min}^{*}}\right).$ (57)
Then, we have
$\left\|(\bm{\Sigma}_{1}\bm{\Sigma}_{1}^{T})^{L-1}\bm{\Sigma}_{1}+\lambda\bm{\Sigma}_{1}-\sqrt{\lambda}\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{A}_{p+1}\right)\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\right\|_{F}\leq c_{2}\|\nabla G(\bm{W})\|_{F},$ (58)
$\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}\leq c_{2}\|\nabla G(\bm{W})\|_{F}.$ (59)
Proof.
Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. It follows from Lemma 2 and 2 that (35)-(37) and (41) holds. Recall that (3.2.1) denotes an SVD of $\bm{W}_{l}$ for each $l\in[L]$. For ease of exposition, let $\bm{H}_{l}:=\bm{U}_{l-1}^{T}\bm{V}_{l}$ and $\bm{T}_{l}:=\mathrm{BlkD}\left(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}\right)$ for each $l=2,\dots,L$. This, together with (48), implies
$\left\|\bm{H}_{l}-\bm{T}_{l}\right\|_{F}\leq\frac{9\sigma_{\max}^{*}}{4\delta_{\sigma}\lambda\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F},\ l=2,\dots,L.$ (60)
First, we are devoted to proving (59). It follows from (41) with $l=L$ that
$\left\|(\bm{W}_{L}\bm{W}^{T}_{L})^{L-1}\bm{W}_{L}-\sqrt{\lambda}\bm{Y}\bm{W}_{L-1:1}^{T}+\lambda\bm{W}_{L}\right\|_{F}\leq c_{1}\|\nabla G(\bm{W})\|_{F},$ (61)
where $c_{1}$ is defined in (42). Using the SVD in (3.2.1), we have
$\textstyle\left\|(\bm{W}_{L}\bm{W}^{T}_{L})^{L-1}\bm{W}_{L}-\sqrt{\lambda}\bm{Y}\bm{W}_{L-1:1}^{T}+\lambda\bm{W}_{L}\right\|_{F}=\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{H}_{l+1}+\lambda\bm{\Sigma}_{L}\right\|_{F}$
$\geq$ $\textstyle\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}-\sqrt{\lambda}\left\|\bm{Y}\bm{V}_{1}\bm{\Sigma}_{1}^{T}(\bm{H}_{2}-\bm{T}_{2})\prod_{l=2}^{L-1}\bm{\Sigma}_{l}^{T}\bm{H}_{l+1}\right\|_{F}$
$\textstyle-\sqrt{\lambda}\left\|\bm{Y}\bm{V}_{1}\bm{\Sigma}_{1}^{T}\bm{T}_{2}\bm{\Sigma}_{2}^{T}(\bm{H}_{3}-\bm{T}_{3})\prod_{l=3}^{L-1}\bm{\Sigma}_{l}^{T}\bm{H}_{l+1}\right\|_{F}-\dots-\sqrt{\lambda}\left\|\bm{Y}\bm{V}_{1}\left(\prod_{l=1}^{L-2}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right)\bm{\Sigma}_{L-1}^{T}(\bm{H}_{L}-\bm{T}_{L})\right\|_{F}$
$\geq$ $\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}-\left(\frac{3}{2}\sigma_{\max}^{*}\right)^{L}\frac{3y_{1}L}{2\sqrt{\lambda}\delta_{\sigma}\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F},$
where the first inequality uses the triangle inequality and the last inequality follows from (35), $\|\bm{H}_{l}\|=\|\bm{T}_{l}\|=1$ for all $l\in[L]$, and (60). This, together with (61), implies
$\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}$ (62)
$\leq$ $\left(\left(\frac{3}{2}\sigma_{\max}^{*}\right)^{L}\frac{3y_{1}L}{2\sqrt{\lambda}\delta_{\sigma}\sigma^{*}_{\min}}+c_{1}\right)\|\nabla G(\bm{W})\|_{F}.$ (63)
Using the diagonal structure of $\bm{\Sigma}_{l}$ in (3.2.1) and $\bm{T}_{l}=\mathrm{BlkD}\left(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}\right)$, we compute
$\textstyle\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}=\mathrm{BlkD}\left(\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(1)}\bm{T}_{l+1}^{(1)},\dots,\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(p)}\bm{T}_{l+1}^{(p)},\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(p+1)}\bm{T}_{l+1}^{(p+1)}\right).$
Using (37) in Lemma 2, we have for each $l\in[L]$ and $i\in[p]$,
$\left\|\bm{\Sigma}_{l}^{(i)}-\bm{\Sigma}_{L}^{(i)}\right\|_{F}^{2}$ $\textstyle=\sum_{j=1}^{g_{i}}\left(\sigma_{j}(\bm{\Sigma}_{l}^{(i)})-\sigma_{j}(\bm{\Sigma}_{L}^{(i)})\right)^{2}=\sum_{j=1}^{g_{i}}\left(\sum_{k=l}^{L-1}\left(\sigma_{j}(\bm{\Sigma}_{k}^{(i)})-\sigma_{j}(\bm{\Sigma}_{k+1}^{(i)})\right)\right)^{2}$
$\leq g_{i}(L-1)^{2}\left(\frac{3\sqrt{2}\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F}\right)^{2}.$
This further implies
$\left\|\bm{\Sigma}_{l}^{(i)}-\bm{\Sigma}_{L}^{(i)}\right\|_{F}\leq\frac{3\sqrt{2g_{\max}}L\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F}.$ (64)
Now we claim that it holds for each $l\in[L]$ and $i\in[p]$ that
$\left\|\bm{T}_{l}^{(i)}\bm{\Sigma}_{L}^{(i)}-\bm{\Sigma}_{L}^{(i)}\bm{T}_{l}^{(i)}\right\|_{F}\leq\frac{\eta_{1}}{\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F},$ (65)
where $\eta_{1}$ is defined in (56). To maintain the flow of the main proof, the proof of this claim is deferred to the end. Using the triangle inequality, we obtain
$\textstyle\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1}\right\|_{F}$
$\leq$ $\ \sum_{j=2}^{L}\|\underbrace{\textstyle\left(\prod_{l=1}^{L-j}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\prod_{k=L-j+2}^{L}\bm{T}_{k}^{(i)}\!\!-\!\!\left(\prod_{k=L-j+2}^{L}\bm{T}_{k}^{(i)}\right)\bm{\Sigma}_{L}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{j-2}}_{\bm{R}_{j}}\!\!\!\!\|_{F}.$ (66)
For each $\|\bm{R}_{j}\|_{F}$, we bound
$\|\bm{R}_{j}\|_{F}$ $\leq\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{L-2}\left\|\bm{\Sigma}_{L}^{(i)}\prod_{k=L-j+2}^{L}\bm{T}_{k}^{(i)}-\left(\prod_{k=L-j+2}^{L}\bm{T}_{k}^{(i)}\right)\bm{\Sigma}_{L}^{(i)}\right\|_{F}$
$\leq\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{L-2}\left\|\left(\bm{\Sigma}_{L}^{(i)}\bm{T}_{L-j+2}^{(i)}-\bm{T}_{L-j+2}^{(i)}\bm{\Sigma}_{L}^{(i)}\right)\prod_{k=L-j+3}^{L}\bm{T}_{k}^{(i)}\right\|_{F}$
$\quad+\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{L-2}\left\|\bm{T}_{L-j+2}^{(i)}\left(\bm{\Sigma}_{L}^{(i)}\bm{T}_{L-j+3}^{(i)}-\bm{T}_{L-j+3}^{(i)}\bm{\Sigma}_{L}^{(i)}\right)\prod_{k=L-j+4}^{L}\bm{T}_{k}^{(i)}\right\|_{F}+\cdots$
$\quad+\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{L-2}\left\|\prod_{k=L-j+2}^{L-1}\bm{T}_{k}^{(i)}(\bm{\Sigma}_{L}^{(i)}\bm{T}_{L}^{(i)}-\bm{T}_{L}^{(i)}\bm{\Sigma}_{L}^{(i)})\right\|_{F}$
[…]
where the first inequality uses (i) of Lemma 2 and $\|\bm{T}_{l+1}^{(i)}\|=1$ for all $l$ and $i$, and the second one is due the triangle inequality. This, together with (3.2.1), implies
$\left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1}\right\|_{F}\leq\left(\frac{3\sigma_{\max}}{2}\right)^{L-2}\frac{L^{2}\eta_{1}}{2\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F}.$ (67)
For each $i\in[p]$, we have
$\textstyle\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1}\right\|_{F}$
$\leq$ $\textstyle\ \left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{L}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)(\bm{\Sigma}_{L}^{(i)})^{L-1}\right\|_{F}+\left\|(\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{L}^{(i)})\bm{T}_{2}^{(i)}\prod_{l=2}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}\right\|_{F}+$
$\textstyle\ \left\|\bm{\Sigma}_{L}^{(i)}\bm{T}_{2}^{(i)}(\bm{\Sigma}_{2}^{(i)}-\bm{\Sigma}_{L}^{(i)})\bm{T}_{3}^{(i)}\prod_{l=3}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}\right\|_{F}+\cdots+\left\|\prod_{l=1}^{L-2}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}(\bm{\Sigma}_{L-1}^{(i)}-\bm{\Sigma}_{L}^{(i)})\bm{T}_{L}^{(i)}\right\|_{F}$
$\leq$ $\ \left(\frac{3\sigma_{\max}^{*}}{2}\right)^{L-2}\left(\frac{L^{2}\eta_{1}}{2\sigma^{*}_{\min}}+\frac{3\sqrt{2g_{\max}}L^{2}\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\right)\|\nabla G(\bm{W})\|_{F},$
where the first inequality uses the triangle inequality, and the last inequality follows from (67), (35) in Lemma 2, $\|\bm{T}_{l}\|=1$ for all $l\in[L]$, and (64). This, together with (55), yields
$\left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}-\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}$
$\leq$ $\ \sum_{i=1}^{p}\left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{(i)}\bm{T}_{l+1}^{(i)}-\left(\prod_{l=1}^{L-1}\bm{T}_{l+1}^{(i)}\right)\left(\bm{\Sigma}_{L}^{(i)}\right)^{L-1}\right\|_{F}$
$\leq$ $\ p\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{L-2}\left(\frac{L^{2}\eta_{1}}{2\sigma^{*}_{\min}}+\frac{3\sqrt{2g_{\max}}L^{2}\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}\right)\|\nabla G(\bm{W})\|_{F}.$ (68)
Now, we are ready to prove (59). Specifically, we have
$\ \left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}$
$\leq$ $\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}\right\|_{F}$
$\quad+\sqrt{\lambda}y_{1}\left\|\prod_{l=1}^{L-1}\bm{\Sigma}_{l}^{T}\bm{T}_{l+1}-\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}$
[…]
where $c_{2}$ is defined in (57). In (61), setting $l=1$ and applying the same argument, we obtain (58).
The rest of the proof is devoted to proving the claim (65). According to (3.2.1) and $\bm{H}_{l}=\bm{U}_{l-1}^{T}\bm{V}_{l}$ for each $l=2,\dots,L$, we have for each $i\in[p]$,
$\left\|\bm{W}_{l}^{T}\bm{W}_{l}-\bm{W}_{l-1}\bm{W}_{l-1}^{T}\right\|_{F}=\left\|\bm{H}_{l}\bm{\Sigma}_{l}^{T}\bm{\Sigma}_{l}\bm{H}_{l}^{T}-\bm{\Sigma}_{l-1}\bm{\Sigma}_{l-1}^{T}\right\|_{F}$
$\geq$ $\left\|\bm{T}_{l}\bm{\Sigma}_{l}^{T}\bm{\Sigma}_{l}\bm{T}_{l}^{T}-\bm{\Sigma}_{l-1}\bm{\Sigma}_{l-1}^{T}\right\|_{F}-2\left\|\bm{\Sigma}_{l}\right\|^{2}\left\|\bm{H}_{l}-\bm{T}_{l}\right\|_{F}$
$\geq$ $\left\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{l}^{(i)})^{2}-(\bm{\Sigma}_{l-1}^{(i)})^{2}\bm{T}_{l}^{(i)}\right\|_{F}-2\left\|\bm{\Sigma}_{l}\|^{2}\|\bm{H}_{l}-\bm{T}_{l}\right\|_{F}$
$\geq$ $\left\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\right\|_{F}-2\left\|\bm{\Sigma}_{l}\right\|^{2}\left\|\bm{H}_{l}-\bm{T}_{l}\right\|_{F}-\left\|(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{l}^{(i)})^{2}\right\|_{F}-\left\|(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{l-1}^{(i)})^{2}\right\|_{F}$
[…]
where the second inequality follows from $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for each $i\in[p]$. This together with (36), (60), and (64), yields that for each $l=2,\dots,L$ and $i\in[p]$, we have
$\left\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\right\|_{F}\leq\eta_{1}\left\|\nabla G(\bm{W})\right\|_{F}.$ (69)
For each $i\in[p]$, using Weyl’s inequality yields $|\sigma_{j}(\bm{\Sigma}_{L}^{(i)})-\sigma_{t_{i}}^{*}|\leq\|\bm{W}_{L}-\bm{W}_{L}^{*}\|\leq{\sigma_{\min}^{*}}/{2}$ for each $j\in[g_{i}]$, which implies $\|\bm{\Sigma}_{L}^{(i)}-\sigma_{t_{i}}^{*}\bm{I}\|\leq\sigma_{\min}^{*}/2$. This, together with Lemma B.4 in Appendix B.2 and $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$, yields $\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\|_{F}\geq\sigma_{\min}^{*}\|\bm{T}_{l}^{(i)}\bm{\Sigma}_{L}^{(i)}-\bm{\Sigma}_{L}^{(i)}\bm{T}_{l}^{(i)}\|_{F}.$ Using this and (69), we obtain (65). ∎
3.2.2 Proof of the Error Bound
Armed with the above results, we now proceed to prove the error bound stated in 2. This is achieved in three steps. First, we show that, in a neighborhood of the critical point set, the singular values and corresponding singular vectors of the weight matrices can be controlled by $\|\nabla G(\bm{W})\|_{F}$ (see Propositions 4 and 5). Second, we establish the error bound for Problem (12) (see 3). To this end, we construct an intermediate point $\hat{\bm{W}}$ using left and right singular matrices (or suitable orthogonal matrices derived from them) of $\bm{W}$ (see (116) in the proof of 3) and show that both $\mathrm{dist}(\hat{\bm{W}},\mathcal{W}_{G})$ and $\|\bm{W}-\hat{\bm{W}}\|_{F}$ are upper bounded by $\|\nabla G(\bm{W})\|_{F}$. Finally, using Lemma 1 and 3, we prove the desired error bound of Problem (1) in 2.
We next bound the singular values and the associated singular vectors of the weight matrices in the following two propositions.
Proposition 4.
Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary. The following statements hold:
(i) Suppose that $L=2$ and (7) holds. Then for all $\bm{W}$ satisfying
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\min\left\{\frac{\delta_{\sigma}}{3},\frac{\sqrt{\lambda}}{\sqrt{3(\sqrt{\lambda}+y_{1})}}\left(\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\sqrt{\lambda}-y_{i}\right|,\sqrt{\lambda}\right\}\right)^{{1}/{2}}\right\},$ (70a)
$\|\nabla G(\bm{W})\|_{F}\leq\frac{\sqrt{2\lambda}\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\sqrt{\lambda}-y_{i}\right|,\sqrt{\lambda}\right\}\sigma^{*}_{\min}}{12c_{2}},$ (70b)
it holds for $l=1,2$ that
$\sigma_{i}(\bm{W}_{l})\leq c_{3}\|\nabla G(\bm{W})\|_{F},\ \forall i=r_{\sigma}+1,\dots,\min\{d_{l},d_{l-1}\},$ (71)
where $c_{3}:={6c_{2}(y_{1}+\sqrt{\lambda})}/({\lambda\min\{\min_{i\in[s_{p_{Y}}]}|\sqrt{\lambda}-y_{i}|,\sqrt{\lambda}\}}).$
(ii) Suppose that $L\geq 3$ and (8) holds. Then for all $\bm{W}$ satisfying
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\min\left\{\frac{\delta_{\sigma}}{3},\left(\frac{\sqrt{\lambda}}{2y_{1}}\right)^{1/(L-2)}\right\},$ (72)
it holds for all $l\in[L]$ that
$\sigma_{i}(\bm{W}_{l})\leq c_{3}\|\nabla G(\bm{W})\|_{F},\ \forall i=r_{\sigma}+1,\dots,\min\{d_{l},d_{l-1}\},$ (73)
where
$c_{3}:=\frac{2}{\lambda}\left(c_{1}+\left(\frac{3}{2}\sigma_{\max}^{*}\right)^{L}\frac{3(L-1)y_{1}}{2\delta_{\sigma}\sqrt{\lambda}\sigma^{*}_{\min}}\right).$ (74)
Proof.
Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. From (70a), we have $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\delta_{\sigma}/3<\sigma_{\min}^{*}/2$ in case (i). From (72), the same bound holds in case (ii). This, together with Lemma 2 and 2, implies that (35)-(37) and (41) hold. According to $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\delta_{\sigma}/3$, 3, and Lemma 3, there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for all $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (48), (58), and (59) hold for all $l=2,\dots,L$, where $\bm{A}_{i}$ and $\bm{B}_{i}$ for each $i\in[p+1]$ are respectively defined in (54) and (55). Recall that (3.2.1) denotes an SVD of $\bm{W}_{l}$ and let $\bm{H}_{l}:=\bm{U}_{l-1}^{T}\bm{V}_{l}$ and $\bm{T}_{l}:=\mathrm{BlkD}(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)})$ for each $l=2,\dots,L$. This, together with (48), implies
$\left\|\bm{H}_{l}-\bm{T}_{l}\right\|_{F}\leq\frac{9\sigma_{\max}^{*}}{4\delta_{\sigma}\lambda\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F},\ l=2,\dots,L.$ (75)
(i) Suppose that $L=2$ and (7) holds. For ease of exposition, define $\bm{\Psi}:=\bm{U}_{2}^{T}\bm{Y}\bm{V}_{1}\in\mathbb{R}^{d_{2}\times d_{0}}$ and we partition $\bm{\Psi}$ as
$\bm{\Psi}=\begin{bmatrix}\bm{\Psi}^{(1,1)}&\cdots&\bm{\Psi}^{(1,p+1)}\\ \vdots&\ddots&\vdots\\ \bm{\Psi}^{(p+1,1)}&\cdots&\bm{\Psi}^{(p+1,p+1)}\end{bmatrix}$, (76)
where $\bm{\Psi}^{(i,j)}\in\mathbb{R}^{g_{i}\times g_{j}}$ for each $i,j\in[p]$, $\bm{\Psi}^{(p+1,j)}\in\mathbb{R}^{(d_{2}-r_{\sigma})\times g_{j}},\ \bm{\Psi}^{(i,p+1)}\in\mathbb{R}^{g_{i}\times(d_{0}-r_{\sigma})}$ for each $i,j\in[p]$, and $\bm{\Psi}^{(p+1,p+1)}\in\mathbb{R}^{(d_{2}-r_{\sigma})\times(d_{0}-r_{\sigma})}$. Using $L=2$, (54), (55), (58) and (59), we have
$c_{2}\|\nabla G(\bm{W})\|_{F}\geq\left\|(\bm{\Sigma}_{1}^{(p+1)}\bm{\Sigma}_{1}^{(p+1)^{T}})\bm{\Sigma}_{1}^{(p+1)}+\lambda\bm{\Sigma}_{1}^{(p+1)}\!\!-\!\!\sqrt{\lambda}\bm{T}_{2}^{(p+1)}\bm{\Sigma}_{2}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}\right\|_{F},$ (77)
$c_{2}\|\nabla G(\bm{W})\|_{F}\geq\left\|(\bm{\Sigma}_{2}^{(p+1)}\bm{\Sigma}_{2}^{(p+1)^{T}})\bm{\Sigma}_{2}^{(p+1)}+\lambda\bm{\Sigma}_{2}^{(p+1)}\!\!-\!\!\sqrt{\lambda}\bm{\Psi}^{(p+1,p+1)}\bm{\Sigma}_{1}^{(p+1)^{T}}\bm{T}_{2}^{(p+1)}\right\|_{F}.$ (78)
Using (58) again, we have
$c_{2}^{2}\|\nabla G(\bm{W})\|_{F}^{2}\geq\left\|(\bm{\Sigma}_{1}\bm{\Sigma}_{1}^{T})\bm{\Sigma}_{1}+\lambda\bm{\Sigma}_{1}-\sqrt{\lambda}\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{A}_{p+1}\right)\bm{\Psi}\right\|_{F}^{2}.$
By dropping the diagonal blocks and the $(p+1)$-th row blocks of the matrix on above the right-hand side, we have
[…]
Applying the same argument to (59) yields $c_{2}^{2}\|\nabla G(\bm{W})\|_{F}^{2}\geq\frac{\lambda\sigma^{*2}_{\min}}{4}\sum_{j=1}^{p}\sum_{i=1,i\neq j}^{p+1}\|\bm{\Psi}^{(i,j)}\|_{F}^{2}.$ This, together with (79), yields
$\frac{8c_{2}^{2}}{\lambda\sigma_{\min}^{*2}}\|\nabla G(\bm{W})\|_{F}^{2}\geq\sum_{i=1}^{p+1}\sum_{j=1,j\neq i}^{p+1}\|\bm{\Psi}^{(i,j)}\|_{F}^{2}.$ (80)
Using (7), we define $\delta:=\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\sqrt{\lambda}-y_{i}\right|,\sqrt{\lambda}\right\}>0.$ Substituting this into (70b) yields
[…]
This, together with $\bm{\Psi}=\bm{U}_{2}^{T}\bm{Y}\bm{V}_{1}$ and Weyl’s inequality, yields for all $i\in[d_{\min}]$,
$\frac{\delta}{3}$ $\geq\left|y_{i}-\sigma_{i}\left(\mathrm{BlkD}(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)})\right)\right|$
$\geq\left|y_{i}-\sqrt{\lambda}\right|-\left|\sigma_{i}\left(\mathrm{BlkD}(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)})\right)-\sqrt{\lambda}\right|$
$\geq\delta-\left|\sigma_{i}\left(\mathrm{BlkD}(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)})\right)-\sqrt{\lambda}\right|,$
which implies $\left|\sigma_{i}(\bm{\Psi}^{(p+1,p+1)})-\sqrt{\lambda}\right|\geq{2\delta}/{3}$ for all $i\in[d_{\min}-r_{\sigma}]$. Accordingly, for all $i\in[d_{\min}-r_{\sigma}]$, we obtain
$\left|\sigma_{i}^{2}(\bm{\Psi}^{(p+1,p+1)})-\lambda\right|$ $=\left|\sigma_{i}(\bm{\Psi}^{(p+1,p+1)})-\sqrt{\lambda}\right|\left|\sigma_{i}(\bm{\Psi}^{(p+1,p+1)})+\sqrt{\lambda}\right|\geq\frac{2\delta\sqrt{\lambda}}{3}.$ (81)
According to $\bm{\Psi}=\bm{U}_{2}^{T}\bm{Y}\bm{V}_{1}$ and (76), we have
$\|\bm{\Psi}^{(p+1,p+1)}\|\leq\|\bm{Y}\|=y_{1}.$ (82)
Using this and (77), we obtain
$c_{2}y_{1}\|\nabla G(\bm{W})\|_{F}$ $\geq\left\|\left((\bm{\Sigma}_{1}^{(p+1)}\bm{\Sigma}_{1}^{(p+1)^{T}})\bm{\Sigma}_{1}^{(p+1)}+\lambda\bm{\Sigma}_{1}^{(p+1)}-\sqrt{\lambda}\bm{T}_{2}^{(p+1)}\bm{\Sigma}_{2}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}\right)\bm{\Psi}^{(p+1,p+1)^{T}}\right\|_{F}$
$\geq\left\|\lambda\bm{\Sigma}_{1}^{(p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}-\sqrt{\lambda}\bm{T}_{2}^{(p+1)}\bm{\Sigma}_{2}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}\right\|_{F}$
$\quad\ -\left\|\bm{\Sigma}_{1}^{(p+1)}\bm{\Sigma}_{1}^{(p+1)^{T}}\bm{\Sigma}_{1}^{(p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}\right\|_{F}$
$\geq\left\|\lambda\bm{T}_{2}^{(p+1)^{T}}\bm{\Sigma}_{1}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)^{T}}-\sqrt{\lambda}\bm{\Sigma}_{2}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}{\bm{\Psi}^{(p+1,p+1)}}^{T}\right\|_{F}-y_{1}\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}^{3},$
where the last inequality follows from $\bm{T}_{2}^{(p+1)}\in\mathcal{O}^{d_{2}-r_{\sigma}}$ and (82). Using (78), we have
$c_{2}\sqrt{\lambda}\|\nabla G(\bm{W})\|_{F}$ $\geq\sqrt{\lambda}\left\|\lambda\bm{\Sigma}_{2}^{(p+1)}-\sqrt{\lambda}\bm{\Psi}^{(p+1,p+1)}\bm{\Sigma}_{1}^{(p+1)^{T}}\bm{T}_{2}^{(p+1)}\right\|_{F}-\sqrt{\lambda}\left\|\bm{\Sigma}_{2}^{(p+1)}\bm{\Sigma}_{2}^{(p+1)^{T}}\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}$
$\geq\left\|\lambda^{\frac{3}{2}}\bm{\Sigma}_{2}^{(p+1)^{T}}-\lambda\bm{T}_{2}^{(p+1)^{T}}\bm{\Sigma}_{1}^{(p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}\right\|_{F}-\sqrt{\lambda}\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}^{3}.$
Summing up the above two inequalities, together with the triangle inequality, yields
$c_{2}(y_{1}+\sqrt{\lambda})\|\nabla G(\bm{W})\|_{F}$ $\geq\left\|\lambda^{\frac{3}{2}}\bm{\Sigma}_{2}^{(p+1)^{T}}-\sqrt{\lambda}\bm{\Sigma}_{2}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}\right\|_{F}$
$\quad-\sqrt{\lambda}\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}^{3}-y_{1}\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}^{3}.$ (83)
Repeating the argument from (82) to (3.2.2), but this time multiplying (77) by $\sqrt{\lambda}$ and (78) by $y_{1}$, we obtain
$c_{2}(y_{1}+\sqrt{\lambda})\|\nabla G(\bm{W})\|_{F}$
$\geq$ $\left\|\lambda^{\frac{3}{2}}\bm{\Sigma}_{1}^{(p+1)^{T}}-\sqrt{\lambda}\bm{\Psi}^{(p+1,p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}\bm{\Sigma}_{1}^{(p+1)^{T}}\right\|_{F}-\sqrt{\lambda}\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}^{3}-y_{1}\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}^{3}.$
Summing up the above two inequalities yields
$2c_{2}(y_{1}+\sqrt{\lambda})\left\|\nabla G(\bm{W})\right\|_{F}$
$\geq$ $\ \sqrt{\lambda}\left\|(\bm{\Psi}^{(p+1,p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}-\lambda\bm{I})\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}+\sqrt{\lambda}\left\|(\bm{\Psi}^{(p+1,p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}-\lambda\bm{I})\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}$
$\quad-(\sqrt{\lambda}+y_{1})\left(\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}^{3}+\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}^{3}\right)$
$\geq$ $\ \sqrt{\lambda}\sigma_{\min}\left(\bm{\Psi}^{(p+1,p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}-\lambda\bm{I}\right)\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}+\sqrt{\lambda}\sigma_{\min}\left(\bm{\Psi}^{(p+1,p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}-\lambda\bm{I}\right)\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}$
$\quad-(\sqrt{\lambda}+y_{1})\left(\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}^{3}+\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}^{3}\right)$
[…]
$\geq$ $\ \frac{\lambda\delta}{3}\left(\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}+\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}\right),$
where the last inequality follows from the definition of $\delta$, (70a), and $\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}=\|\bm{\Sigma}_{1}^{(p+1)}-\bm{0}\|_{F}\leq\|\bm{\Sigma}_{1}-\bm{\Sigma}_{1}^{*}\|_{F}\leq\|\bm{W}_{1}-\bm{W}_{1}^{*}\|_{F}\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})$ and $\|\bm{\Sigma}_{2}^{(p+1)}\|_{F}\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})$ due to Mirsky’s inequality (see Lemma D.8 in Appendix D). This directly implies (71).
(ii) Suppose that $L\geq 3$ and (8) holds. For ease of exposition, let $\bm{\Psi}:=\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\in\mathbb{R}^{d_{L}\times d_{0}}$ and we still partition $\bm{\Psi}$ into the form as in (76). Let $l\in[L]$ be such that $l\in\operatorname{\arg\max}\{\sigma_{r_{\sigma}+1}(\bm{W}_{k}):{k\in[L]}\}$. Define
$\bm{\Delta}$ $\textstyle:=\left(\prod_{k=l+1}^{L}\bm{H}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right)-\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)$
$\textstyle~=\sum_{j=l}^{L-1}\left(\prod_{k=l+1}^{j}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)(\bm{H}_{j+1}-\bm{T}_{j+1})\bm{\Sigma}_{j+1}^{T}\left(\prod_{k=j+2}^{L}\bm{H}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right)$
$\textstyle\quad+\sum_{j=2}^{l}\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{j-2}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)\bm{\Sigma}_{j-1}^{T}(\bm{H}_{j}-\bm{T}_{j})\left(\prod_{k=j}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right).$
Here, we remark that in the above summations, if the lower index exceeds the upper index, the corresponding product is interpreted as the identity matrix. Then substituting (3.2.1) into (41) yields
$c_{1}\|\nabla G(\bm{W})\|_{F}$ $\textstyle\geq\left\|(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{H}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right)+\lambda\bm{\Sigma}_{l}\right\|_{F}$
$\textstyle\geq\left\|(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)+\lambda\bm{\Sigma}_{l}\right\|_{F}-\sqrt{\lambda}\|\bm{\Delta}\|_{F},$
where the second inequality uses the triangle inequality. Using (35) and (75), we obtain
$\|\bm{\Delta}\|\leq\left(\frac{3}{2}\sigma_{\max}^{*}\right)^{L}\frac{3(L-1)y_{1}}{2\delta_{\sigma}\lambda\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F}.$
This, together with the above inequality, yields
$\textstyle\left\|(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)+\lambda\bm{\Sigma}_{l}\right\|_{F}\leq\eta_{2}\|\nabla G(\bm{W})\|_{F},$ (84)
where $\eta_{2}:=c_{1}+\left(\frac{3}{2}\sigma_{\max}^{*}\right)^{L}\frac{3(L-1)y_{1}}{2\delta_{\sigma}\sqrt{\lambda}\sigma^{*}_{\min}}$. Using this inequality, the block structures of $\bm{\Sigma}_{l}$ and $\bm{T}_{l}$, $l\in\operatorname{\arg\max}\{\sigma_{r_{\sigma}+1}(\bm{W}_{k}):{k\in[L]}\}$, and the fact that the Frobenius norm of a matrix is larger than or equal to its spectral norm, we obtain
$\eta_{2}\|\nabla G(\bm{W})\|_{F}$
$\geq$ $\left\|\left(\bm{\Sigma}_{l}^{(p+1)}\bm{\Sigma}_{l}^{(p+1)^{T}}\right)^{L-1}\bm{\Sigma}_{l}^{(p+1)}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{T}_{k}^{(p+1)}\bm{\Sigma}_{k}^{(p+1)^{T}}\right)\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{(p+1)^{T}}\bm{T}_{k+1}^{(p+1)}\right)+\lambda\bm{\Sigma}_{l}^{(p+1)}\right\|$
$\geq$ $\ \left\|\left(\bm{\Sigma}_{l}^{(p+1)}\bm{\Sigma}_{l}^{(p+1)^{T}}\right)^{L-1}\bm{\Sigma}_{l}^{(p+1)}+\lambda\bm{\Sigma}_{l}^{(p+1)}\right\|-\sqrt{\lambda}\sigma_{r_{\sigma}+1}^{L-1}(\bm{W}_{l})\left\|\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\right\|$
$\geq$ $\ \lambda\sigma_{r_{\sigma}+1}(\bm{W}_{l})-\sqrt{\lambda}y_{1}\sigma_{r_{\sigma}+1}^{L-1}(\bm{W}_{l})\geq\sqrt{\lambda}\left(\sqrt{\lambda}-y_{1}\sigma_{r_{\sigma}+1}^{L-2}(\bm{W}_{l})\right)\sigma_{r_{\sigma}+1}(\bm{W}_{l}),$
where $\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}$ denotes the submatrix of $\bm{\Psi}$ with the rows indexed from $r_{\sigma}+1$ to $d_{L}$ and the columns indexed from $r_{\sigma}+1$ to $d_{0}$, and the third inequality is due to $\|\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\|\leq\|\bm{Y}\|=y_{1}$. Using Weyl’s inequality and (72), we have $\sigma_{r_{\sigma}+1}(\bm{W}_{l})\leq\mathrm{dist}\left(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}}\right)\leq\left(\frac{\sqrt{\lambda}}{2y_{1}}\right)^{1/(L-2)}.$ This, together with the above inequality , directly yields $\sigma_{r_{\sigma}+1}(\bm{W}_{l})\leq{2\eta_{2}}\|\nabla G(\bm{W})\|_{F}/{\lambda}$, which implies (73). ∎
The above proposition provides bounds for the $\min\{d_{l},d_{l-1}\}-r_{\sigma}$ smallest singular values of $\bm{W}_{l}$. It remains to bound the top $r_{\sigma}$ singular values and all singular vectors. For convenience, we define an auxiliary function $\varphi:\mathbb{R}_{++}\;\to\;\mathbb{R}$:
$\varphi(x):=\frac{x^{2L-1}+\lambda x}{\sqrt{\lambda}\,x^{L-1}}.$ (85)
Proposition 5.
Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary and $\bm{W}$ be arbitrary such that
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\min\left\{\delta_{1},\delta_{2},\frac{\delta_{\sigma}}{3}\right\}\quad\text{and}\quad\|\nabla G(\bm{W})\|_{F}\leq\frac{\delta_{y}\sqrt{\lambda}(\sigma^{*}_{\min})^{L-1}}{3\cdot 2^{L-1}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}},$ (86)
where
$\delta_{1}$ $:=\frac{\delta_{y}}{{3L}\left(\frac{4\sigma_{\max}^{*}}{3}\right)^{L-1}\frac{1}{\sqrt{\lambda}}+3(L-2)\sqrt{\lambda}\left(\frac{2\sigma_{\min}^{*}}{3}\right)^{1-L}},$
$\delta_{2}$ $:=\frac{\min_{i\in[r_{\sigma}]}|\varphi^{\prime}(\sigma_{i}^{*})|}{{2L(L-1)}\left(\frac{4\sigma_{\max}^{*}}{3}\right)^{L-1}\frac{1}{\sqrt{\lambda}}+2\sqrt{\lambda}(2-L)(1-L)\left(\frac{2\sigma_{\min}^{*}}{3}\right)^{-L}},$ (87)
$\ \eta_{3}$ $:=c_{2}+c_{3}\sqrt{d_{\max}}\left(\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(L-1)}\!\!\!\!+\sqrt{\lambda}y_{1}\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{L-2}\!\!\!\!+\lambda\right),$
$\eta_{4}$ $:=\eta_{3}+\frac{p\eta_{1}(2L-1)L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}\!\!\!\!+\frac{\lambda p\eta_{1}L}{\sigma^{*}_{\min}}.$ (88)
Suppose in addition Assumptions 1 and 2 hold, (70) holds for $L=2$, and (72) holds for $L\geq 3$.
(i) There exist orthogonal matrices $\hat{\bm{U}}_{L}^{(i)}\in\mathcal{O}^{h_{i}}$ for $i\in[p_{Y}]$, $\hat{\bm{U}}_{L}^{(p_{Y}+1)}\in\mathcal{O}^{d_{L}-r_{Y}}$, $\hat{\bm{V}}_{L}^{(p_{Y}+1)}\in\mathcal{O}^{d_{0}-r_{Y}}$, $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for $i\in[p]$, $\bm{P}\in\mathcal{O}^{d_{L}-r_{\sigma}}$, $\bm{Q}\in\mathcal{O}^{d_{0}-r_{\sigma}}$, and a permutation matrix $\bm{\Pi}\in\mathcal{P}^{d_{\min}}$ with $(\bm{\sigma}^{*},\bm{\Pi}^{T})\in\mathcal{B}$ such that
$\|\tilde{\bm{U}}_{L}-\bm{U}_{L}\|_{F}\leq c_{4}\|\nabla G(\bm{W})\|_{F}\quad\text{ and }\quad\|\tilde{\bm{V}}_{1}-\bm{V}_{1}\|_{F}\leq c_{4}\|\nabla G(\bm{W})\|_{F},$ (89)
where
$\eta_{5}:=\frac{2^{L+1}(6y_{1}+\delta_{y})(p_{Y}+1)\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{3\sqrt{\lambda}y_{s_{p_{Y}}}\delta_{y}\sigma^{*L-1}_{\min}},\ c_{4}:=\eta_{5}+\frac{1}{y_{s_{p_{Y}}}}\left(\frac{2^{L}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma^{*L-1}_{\min}}+2y_{1}\eta_{5}\right)$ (90)
$\tilde{\bm{U}}_{L}:=\mathrm{BlkD}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\mathrm{BlkD}\left(\bm{\Pi},\bm{I}_{d_{L}-d_{\min}}\right)\mathrm{BlkD}\left(\bm{I}_{r_{\sigma}},\bm{P}^{T}\right),$ (91)
$\tilde{\bm{V}}_{1}:=\mathrm{BlkD}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{V}}_{L}^{(p_{Y}+1)}\right)\mathrm{BlkD}\left(\bm{\Pi},\bm{I}_{d_{0}-d_{\min}}\right)\mathrm{BlkD}\left(\bm{I}_{r_{\sigma}},\bm{Q}^{T}\right)\hat{\bm{T}}$, (92)
$\hat{\bm{T}}:=\mathrm{BlkDiag}\left(\prod_{l=2}^{L}\bm{T}_{l}^{(1)},\cdots,\prod_{l=2}^{L}\bm{T}_{l}^{(p)},\bm{I}_{d_{0}-r_{\sigma}}\right).$ (93)
(ii) For each $l\in[L]$, there exists a constant $c_{5}>0$ such that
$|\sigma_{i}\left(\bm{W}_{l})-\sigma^{*}_{i}\right|\leq c_{5}\|\nabla G(\bm{W})\|_{F},\ \forall i\in[r_{\sigma}],$ (94)
where
$c_{5}:=\frac{2^{L}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma_{\min}^{*{L-1}}\min_{i\in[r_{\sigma}]}\left|\varphi^{\prime}(\sigma_{i}^{*})\right|}+\frac{3\sqrt{2}(L-1)\sigma_{\max}^{*}}{4\lambda\sigma_{\min}^{*}}.$ (95)
Proof.
According to (70a) and (72), we have $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\delta_{\sigma}/3<\sigma_{\min}^{*}/2$. This, together with Lemma 2 and 2, implies that (35)-(37) and (41) hold. According to 3 and Lemma 3 with $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\delta_{\sigma}/3$, there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for each $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (58) and (59) hold, where $\bm{A}_{i}$ and $\bm{B}_{i}$ for $i\in[p+1]$ are defined in (54) and (55). Based on (70) (resp., (72)) and 4, we have (71) (resp., (73)). Recall that (3.2.1) denotes an SVD of $\bm{W}_{l}$ for $l\in[L]$. Let
$\bm{\Psi}:=\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1},\qquad\hat{\bm{\Sigma}}_{l}:=\mathrm{BlkD}\left(\bm{\Sigma}_{l}^{(1)},\dots,\bm{\Sigma}_{l}^{(p)},\bm{0}_{(d_{l}-r_{\sigma})\times(d_{l-1}-r_{\sigma})}\right),\ \forall l\in[L].$ (96)
(i) We first prove the first inequality in (89). We claim that the following two inequalities hold:
$\sum_{i=1}^{r_{\sigma}}\left\|\varphi(\sigma_{i}(\bm{W}_{L}))\bm{e}_{i}-\bm{\Psi}\hat{\bm{T}}\bm{e}_{i}\right\|^{2}$ $\leq\frac{4^{L-1}\eta_{3}^{2}}{\lambda(\sigma_{\min}^{*})^{2L-2}}\|\nabla G(\bm{W})\|_{F}^{2},$ (97)
$\sum_{i=1}^{r_{\sigma}}\left\|\varphi(\sigma_{i}(\bm{W}_{1}))\bm{e}_{i}^{T}-\bm{e}_{i}^{T}\bm{\Psi}\hat{\bm{T}}\right\|^{2}$ $\leq\frac{4^{L-1}\eta_{4}^{2}}{\lambda(\sigma_{\min}^{*})^{{2L-2}}}\|\nabla G(\bm{W})\|_{F}^{2}.$ (98)
To maintain the flow of the main proof, the proof of this claim is deferred to Appendix B.3.1. Now, let $\bm{\alpha}:=\left(\varphi(\sigma_{1}(\bm{W}_{L})),\dots,\varphi(\sigma_{r_{\sigma}}(\bm{W}_{L}))\right)\in\mathbb{R}^{r_{\sigma}}$ and $\bm{Z}:=\mathrm{BlkDiag}\left(\mathrm{diag}(\bm{\alpha}),(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\right)$. We have
$\left\|\bm{\Psi}\hat{\bm{T}}-\bm{Z}\right\|_{F}^{2}$ $=\sum_{i=1}^{r_{\sigma}}\left\|(\bm{\Psi}\hat{\bm{T}}-\bm{Z})\bm{e}_{i}\right\|^{2}+\sum_{i=r_{\sigma}+1}^{d_{0}}\|(\bm{\Psi}\hat{\bm{T}}-\bm{Z})\bm{e}_{i}\|^{2}$
$=\sum_{i=1}^{r_{\sigma}}\left\|\bm{\Psi}\hat{\bm{T}}\bm{e}_{i}-\varphi(\sigma_{i}(\bm{W}_{L}))\bm{e}_{i}\right\|^{2}+\left\|\left(\bm{\Psi}\hat{\bm{T}}\right)_{1:r_{\sigma},r_{\sigma}+1:d_{0}}\right\|_{F}^{2}$
$\leq\sum_{i=1}^{r_{\sigma}}\left\|\bm{\Psi}\hat{\bm{T}}\bm{e}_{i}-\varphi(\sigma_{i}(\bm{W}_{L}))\bm{e}_{i}\right\|^{2}+\sum_{i=1}^{r_{\sigma}}\left\|\bm{e}_{i}^{T}\bm{\Psi}\hat{\bm{T}}-\varphi(\sigma_{i}(\bm{W}_{1}))\bm{e}_{i}^{T}\right\|^{2}$
[…]
where the second equality uses the structure of $\bm{Z}$ and the first inequality is due to
$\left\|\bm{e}_{i}^{T}\bm{\Psi}\hat{\bm{T}}-\varphi(\sigma_{i}(\bm{W}_{1}))\bm{e}_{i}^{T}\right\|\geq\|(\bm{\Psi}\hat{\bm{T}})_{i,r_{\sigma}+1:d_{0}}\|\ \text{for all}\ \ i\in[r_{\sigma}].$
Let $(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}=\bm{P}\bm{\Lambda}\bm{Q}^{T}$ be an SVD of $(\bm{\Psi}\hat{\bm{T}})_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}$, where $\bm{\Lambda}\in\mathbb{R}^{(d_{L}-r_{\sigma})\times(d_{0}-r_{\sigma})}$ is diagonal with entries $\gamma_{1}\geq\dots\geq\gamma_{d_{\min}-r_{\sigma}}$ being the top $d_{\min}-r_{\sigma}$ singular values. To proceed, we define
$\bm{\beta}:=\left(\varphi(\sigma_{1}(\bm{W}_{L})),\dots,\varphi(\sigma_{r_{\sigma}}(\bm{W}_{L})),\gamma_{1},\dots,\gamma_{d_{\min}-r_{\sigma}}\right)\in\mathbb{R}^{d_{\min}}.$ (100)
Let $\bm{\Pi}\in\mathcal{P}^{d_{\min}}$ be a permutation matrix such that the entries $\bm{\Pi}\bm{\beta}$ are in a decreasing order and
$\bar{\bm{U}}_{L}=\bm{U}_{L}\mathrm{BlkDiag}\left(\bm{I}_{r_{\sigma}},\bm{P}\right)\mathrm{BlkDiag}\left(\bm{\Pi}^{T},\bm{I}_{d_{L}-d_{\min}}\right),$ (101)
$\bar{\bm{V}}_{1}=\bm{V}_{1}\hat{\bm{T}}\mathrm{BlkDiag}\left(\bm{I}_{r_{\sigma}},\bm{Q}\right)\mathrm{BlkDiag}\left(\bm{\Pi}^{T},\bm{I}_{d_{0}-d_{\min}}\right),$
$\bar{\bm{Y}}=\mathrm{BlkDiag}\left(\bm{\Pi}\mathrm{diag}\left(\bm{\beta}\right)\bm{\Pi}^{T},\bm{0}_{(d_{L}-d_{\min})\times(d_{0}-d_{\min})}\right).$
Since both $\bm{Y}$ and $\bar{\bm{Y}}$ are diagonal matrices with elements in decreasing order, by using Mirsky’s inequality (see Lemma D.8 in Appendix D), we have
[…]
where the equality follows from (140), (100), and (101). This, together with (86), implies
$\|\bm{Y}-\bar{\bm{Y}}\|_{F}\leq\frac{\delta_{y}}{3},\qquad\|\bar{\bm{Y}}\|\leq\|\bar{\bm{Y}}-\bm{Y}\|+\|\bm{Y}\|\leq y_{1}+\frac{\delta_{y}}{3}.$ (103)
Using the triangle inequality, we obtain
$\quad\quad\|\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}-\bar{\bm{Y}}\bar{\bm{Y}}^{T}\|_{F}$
$~~~\leq\|(\bar{\bm{U}}_{L}^{T}\bm{Y}\bar{\bm{V}}_{1}-\bar{\bm{Y}})\bar{\bm{V}_{1}}^{T}\bm{Y}^{T}\bar{\bm{U}}_{L}\|_{F}+\|\bar{\bm{Y}}(\bar{\bm{U}}_{L}^{T}\bm{Y}\bar{\bm{V}}_{1}-\bar{\bm{Y}})^{T}\|_{F}$
[…]
Next, we write $\bar{\bm{U}}_{L}$ and $\bm{I}_{d_{L}}$ as follows:
$\bar{\bm{U}}_{L}=\begin{bmatrix}(\bar{\bm{U}}_{L}^{(1)})^{T}&\cdots&(\bar{\bm{U}}_{L}^{(p_{Y})})^{T}&(\bar{\bm{U}}_{L}^{(p_{Y}+1)})^{T}\end{bmatrix}^{\!T},\ \bm{I}_{d_{L}}=\begin{bmatrix}(\bm{E}^{(1)})^{T}&\cdots&(\bm{E}^{(p_{Y})})^{T}&(\bm{E}^{(p_{Y}+1)})^{T}\end{bmatrix}^{\!T},$
$\bm{E}^{(1)}=[\,\bm{I}_{h_{1}}\ \ \bm{0}_{h_{1}\times(d_{L}-s_{1})}\,],\ \bm{E}^{(i)}=[\,\bm{0}_{h_{i}\times s_{i-1}}\ \ \bm{I}_{h_{i}}\ \ \bm{0}_{h_{i}\times(d_{L}-s_{i})}\,],\ \bm{E}^{(p_{Y}+1)}=[\,\bm{0}_{(d_{L}-r_{Y})\times r_{Y}}\ \ \bm{I}_{d_{L}-r_{Y}}\,],$
where $\bar{\bm{U}}_{L}^{(i)},\ \bm{E}^{(i)}\in\mathbb{R}^{h_{i}\times d_{L}}$ for all $i\in[p_{Y}]$, and $\bar{\bm{U}}_{L}^{(p_{Y}+1)},\ \bm{E}^{(p_{Y}+1)}\in\mathbb{R}^{(d_{L}-r_{Y})\times d_{L}}$. Applying the Davis-Kahan theorem (see Lemma D.7 in Appendix D) to the matrices $\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}$ (considered as $\bm{A}$), $\bar{\bm{Y}}\bar{\bm{Y}}^{T}$ (considered as $\hat{\bm{A}}$), $\bar{\bm{U}}_{L}^{(i)^{T}}$ (considered as $\bm{V}$) , and $\bm{E}^{(i)^{T}}$ (considered as $\hat{\bm{V}}$) for each $i\in[p_{Y}+1]$, there exist orthogonal matrices $\hat{\bm{U}}_{L}^{(i)}\in\mathcal{O}^{h_{i}}$ for each $i\in[p_{Y}]$ and $\hat{\bm{U}}_{L}^{(p_{Y}+1)}\in\mathcal{O}^{d_{L}-r_{Y}}$ such that
$\left\|\bar{\bm{U}}_{L}^{(i)}-\hat{\bm{U}}_{L}^{(i)}\bm{E}^{(i)}\right\|_{F}\leq\frac{4\|\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}-\bar{\bm{Y}}\bar{\bm{Y}}^{T}\|_{F}}{\min\{\lambda_{i-1}-\lambda_{i},\lambda_{i}-\lambda_{i+1}\}},\ \forall i\in[p_{Y}+1],$ (105)
where $\lambda_{i}$ is the $i$-th largest non-repeated eigenvalue of $\bar{\bm{U}}_{L}^{T}\bm{Y}\bm{Y}^{T}\bar{\bm{U}}_{L}$ and $\lambda_{0}=\infty$ and $\lambda_{p_{Y}+2}=$-$\infty$ by convention. We compute $\lambda_{i}=y_{s_{i}}^{2}$ for $i\in[p_{Y}]$ and $\lambda_{p_{Y}+1}=0$ for $i\in[p_{Y}]$. Using (104) and the following inequalities
$y_{s_{j}}^{2}-y_{s_{j}+1}^{2}=(y_{s_{j}}-y_{s_{j}+1})(y_{s_{j}}+y_{s_{j}+1})\geq\delta_{y}y_{s_{p_{Y}}},\ \forall j\in[p_{Y}],$
$\min\left\{y_{s_{i}}^{2}-y_{s_{i}+1}^{2},y_{s_{i+1}}^{2}-y_{s_{i+1}+1}^{2}\right\}\geq\delta_{y}y_{s_{p_{Y}}},\ \forall i\in[p_{Y}-1],$
we have from (105) that $\|\bar{\bm{U}}_{L}^{(i)}-\hat{\bm{U}}_{L}^{(i)}\bm{E}^{(i)}\|_{F}\leq\frac{2^{L+1}(6y_{1}+\delta_{y})\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{3\sqrt{\lambda}y_{s_{p_{Y}}}\delta_{y}\sigma^{*L-1}_{\min}}\|\nabla G(\bm{W})\|_{F}.$ Therefore, we have
$\left\|\bar{\bm{U}}_{L}\!-\!\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\right\|_{F}$ $\!\leq\!\sum_{i=1}^{p_{Y}+1}\left\|\bar{\bm{U}}_{L}^{(i)}-\hat{\bm{U}}_{L}^{(i)}\bm{E}^{(i)}\right\|_{F}\leq\eta_{5}\|\nabla G(\bm{W})\|_{F}.$ (106)
According to definition of $\tilde{\bm{U}}_{L}$ in (91), we have
[…]
Next we prove the second inequality in (89). Using the same argument in (104)-(106) to $\bar{\bm{V}}_{1}^{T}\bm{Y}^{T}\bm{Y}\bar{\bm{V}}_{1}$ and $\bar{\bm{Y}}^{T}\bar{\bm{Y}}$, we conclude that there exist orthogonal matrices $\hat{\bm{V}}_{1}^{(i)}\in\mathcal{O}^{h_{i}}$ for all $i\in[p]$ and $\hat{\bm{V}}_{1}^{(p+1)}\in\mathcal{O}^{d_{0}-r_{Y}}$ such that
$\left\|\bar{\bm{V}}_{1}-\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)\right\|_{F}\leq\eta_{5}\|\nabla G(\bm{W})\|_{F}.$ (107)
Using the block structure of $\bm{Y}$ in (4), we have
$\ \left\|\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})}\right)-\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})}\right)\right\|_{F}$
$\leq$ $\ \frac{1}{y_{s_{p_{Y}}}}\left\|\mathrm{BlkDiag}\left(y_{s_{1}}\hat{\bm{U}}_{L,1}\hat{\bm{V}}_{1,1}^{T},\dots,y_{s_{p_{Y}}}\hat{\bm{U}}_{L,p_{Y}}\hat{\bm{V}}_{1,p_{Y}}^{T}\right)-\mathrm{BlkDiag}\left(y_{s_{1}}\bm{I}_{h_{1}},\dots,y_{s_{p_{Y}}}\bm{I}_{h_{p_{Y}}}\right)\right\|_{F}$
$=$ $\ \frac{1}{y_{s_{p_{Y}}}}\left\|\mathrm{BlkD}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)\bm{Y}\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)^{T}-\bm{Y}\right\|_{F}$
$\leq$ $\ \frac{1}{y_{s_{p_{Y}}}}\left(\left\|\left(\mathrm{BlkD}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{U}}_{L}^{(p_{Y}+1)}\right)-\bar{\bm{U}}_{L}\right)\bm{Y}\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)^{T}\right\|_{F}\right$.
$\hskip 18.49988pt~~~\left.+\left\|\bar{\bm{U}}_{L}\bm{Y}\left(\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)-\bar{\bm{V}}_{1}\right)^{T}\right\|_{F}+\|\bar{\bm{U}}_{L}\bm{Y}\bar{\bm{V}}_{1}^{T}-\bar{\bm{Y}}\|_{F}+\|\bar{\bm{Y}}-\bm{Y}\|_{F}\right)$
$\leq$ $\ \frac{1}{y_{s_{p_{Y}}}}\left(\frac{2^{L}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma^{*L-1}_{\min}}+2y_{1}\eta_{5}\right)\|\nabla G(\bm{W})\|_{F}.$
where the last inequality is due to (102), (106), and (107). Using the definition of $\tilde{\bm{V}}_{1}$, we compute
$~~~~\|\bm{V}_{1}-\tilde{\bm{V}}_{1}\|_{F}=\left\|\bar{\bm{V}}_{1}-\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)\right\|_{F}$
$\leq\left\|\bar{\bm{V}}_{1}-\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})},\hat{\bm{V}}_{1}^{(p_{Y}+1)}\right)\right\|_{F}$
$\quad+\left\|\mathrm{BlkD}\left(\hat{\bm{V}}_{1}^{(1)},\dots,\hat{\bm{V}}_{1}^{(p_{Y})}\right)-\mathrm{BlkDiag}\left(\hat{\bm{U}}_{L}^{(1)},\dots,\hat{\bm{U}}_{L}^{(p_{Y})}\right)\right\|_{F}$
[…]
That is, the second inequality in (89) holds.
The rest of the proof is devoted to showing that $\bm{\Pi}$ satisfies $(\bm{\sigma}^{*},\bm{\Pi}^{T})\in\cal$ B. According to (103) and the definition of $\bar{\bm{Y}}$ in (101), we have $\left\|\bm{\beta}-\bm{\Pi}^{T}\bm{y}\right\|\leq{\delta_{y}}/{3},$ where $\bm{y}:=(y_{1},y_{2},\dots,y_{d_{\min}})$. This implies that there exists a permutation $\pi:[d_{\min}]\to[d_{\min}]$ such that
$\left|\beta_{i}-y_{\pi^{-1}(i)}\right|\leq\frac{\delta_{y}}{3},\ \forall i\in[d_{\min}].$ (108)
Using Weyl’s inequality and (86), we have
$\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\min\left\{\delta_{1},\frac{\delta_{\sigma}}{3}\right\},$
This, together with (100) and Lemma B.5 in Appendix B.3, implies
$\left|\beta_{i}-\varphi(\sigma_{i}^{*})\right|=\left|\varphi(\sigma_{i}(\bm{W}_{L}))-\varphi(\sigma_{i}^{*})\right|\leq\frac{\delta_{y}}{3},\ \forall i\in[r_{\sigma}].$ (109)
According to (26), there exists $k_{i}\in[d_{\min}]$ such that $\varphi(\sigma_{i}^{*})=y_{k_{i}}$ for all $i\in[r_{\sigma}]$. For all $i\in[r_{\sigma}]$, we have
[…]
Combining this with the definition of $\delta_{y}$ in (18) yields $\varphi(\sigma_{i}^{*})=y_{k_{i}}=y_{\pi^{-1}(i)}$ for all $i\in[r_{\sigma}]$. Thus, we have
$(\sigma^{*}_{i})^{2L-1}+\lambda\sigma^{*}_{i}-\sqrt{\lambda}y_{\pi^{-1}(i)}(\sigma^{*}_{i})^{{L-1}}=0,\ \forall i\in[r_{\sigma}].$
For each $i\in\{r_{\sigma}+1,\dots,d_{\min}\}$, we have $\sigma^{*}_{i}=0$, and it is trivial to see that the above equation holds. Thus, we have $(\sigma_{\pi(i)}^{*})^{2L-1}-\sqrt{\lambda}\,y_{i}\,(\sigma_{\pi(i)}^{*})^{L-1}+\lambda\,\sigma^{*}_{\pi(i)}=0,\ \sigma_{\pi(i)}^{*}\geq 0,\ \forall i\in[d_{\min}].$ This, together with the definition of $\mathcal{A}$ in (26), yields $\bm{a}:=\bm{\Pi}\bm{\sigma}^{*}=\left(\sigma_{\pi(1)}^{*},\ldots,\sigma_{\pi(d_{\min})}^{*}\right)\in\mathcal{A}$. It follows from $\bm{\sigma}^{*}=\bm{\Pi}^{T}\bm{a}$ and the definition of $\mathcal{B}$ in (27) that $(\bm{\sigma}^{*},\bm{\Pi}^{T})\in\mathcal{B}$.
(ii) Using (100) and the fact that $y_{\pi^{-1}(i)}=\varphi(\sigma_{i}^{*})$ for each $i\in[r_{\sigma}]$, we have
$\left|\varphi(\sigma_{i}(\bm{W}_{L}))-\varphi(\sigma_{i}^{*})\right|$ $=\left|\beta_{i}-y_{\pi^{-1}(i)}\right|\leq\|\bm{\beta}-\bm{\Pi}^{T}\bm{y}\|=\|\bm{\Pi}\bm{\beta}-\bm{y}\|$
[…]
Under Assumption 2, by Lemma A.2 in Appendix A.2, we have $\varphi^{\prime}(\sigma_{i}^{*})\neq 0$ for each $i\in[r_{\sigma}]$, so that $\min_{i\in[r_{\sigma}]}\left|\varphi^{\prime}(\sigma_{i}^{*})\right|>0$. This, together with Lemma B.5 (see Appendxi B.3), yields that for all $i\in[r_{\sigma}]$,
$\left|\varphi^{\prime}(x)-\varphi^{\prime}(\sigma_{i}^{*})\right|\leq\frac{|\varphi^{\prime}(\sigma_{i}^{*})|}{2},\quad\text{if }|x-\sigma_{i}^{*}|\leq\min\left\{\delta_{1},\delta_{2},\frac{\delta_{\sigma}}{3}\right\}.$ (111)
By Weyl’s inequality, we have
[…]
Applying the mean value theorem to $\varphi(\cdot)$, there exists $x$ satisfying $\sigma_{i}(\bm{W}_{L})\leq x\leq\sigma_{i}^{*}$ such that
$\left|\varphi(\sigma_{i}(\bm{W}_{L}))-\varphi(\sigma_{i}^{*})\right|=\left|\varphi^{\prime}(x)\right|\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\geq\frac{\left|\varphi^{\prime}(\sigma_{i}^{*})\right|}{2}\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|,$
where the inequality follows from (111) and $|x-\sigma_{i}^{*}|\leq\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\leq\delta_{2}$ due to (112). This, together with (3.2.2), yields for all $i\in[r_{\sigma}]$,
$\left|\sigma_{i}(\bm{W}_{L})-\sigma_{i}^{*}\right|\leq\frac{2^{L}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}}{\sqrt{\lambda}\sigma_{\min}^{*{L-1}}\min_{i\in[r_{\sigma}]}\left|\varphi^{\prime}(\sigma_{i}^{*})\right|}\|\nabla G(\bm{W})\|_{F}.$
Using this and (37) in Lemma 2, we obtain the desired ∎
Now we are ready to prove the error bound for Problem (12).
Theorem 3.
Suppose that Assumptions 1 and  2 hold. Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary and $\bm{W}$ be arbitrary such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\epsilon_{\bm{\sigma}^{*}}$. Then we have
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\kappa_{\bm{\sigma}^{*}}\|\nabla G(\bm{W})\|_{F},$ (113)
where
$\epsilon_{\bm{\sigma}^{*}}$ $:=\begin{cases}\min\!\left\{\begin{array}[]{@{}l@{}}\frac{\delta_{\sigma}}{3},\;\delta_{1},\;\delta_{2},\;\frac{\delta_{y}\sqrt{\lambda}\,(\sigma^{*}_{\min})^{L-1}}{3\cdot 2^{L-1}L_{G}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}},\\[4.30554pt] \quad\frac{\sqrt{\lambda}}{\sqrt{3(\sqrt{\lambda}+y_{1})}}\Biggl(\min\!\bigl\{\,\min_{i\in[s_{p_{Y}}]}\lvert\sqrt{\lambda}-y_{i}\rvert,\;\sqrt{\lambda}\,\bigr\}\Biggr)^{\!1/2},\\[5.16663pt] \quad\frac{\sqrt{2\lambda}\,\min\!\bigl\{\,\min_{i\in[s_{p_{Y}}]}\lvert\sqrt{\lambda}-y_{i}\rvert,\;\sqrt{\lambda}\,\bigr\}\,\sigma^{*}_{\min}}{12\,c_{2}\,L_{G}}\end{array}\right\},&\text{if}\ L=2,\\[43.05542pt] \min\!\Biggl\{\frac{\delta_{\sigma}}{3},\;\delta_{1},\;\delta_{2},\;\frac{\delta_{y}\sqrt{\lambda}\,(\sigma^{*}_{\min})^{L-1}}{3\cdot 2^{L-1}L_{G}\sqrt{\eta_{3}^{2}+\eta_{4}^{2}}},\;\left(\frac{\sqrt{\lambda}}{2y_{1}}\right)^{\!\frac{1}{L-2}}\Biggr\},&\text{if}\ L\geq 3,\end{cases}$ (114)
$\kappa_{\bm{\sigma}^{*}}$ $:=\sqrt{L}\left(\frac{9\sigma_{\max}^{*2}}{4\delta_{\sigma}\lambda\sigma^{*}_{\min}}+c_{3}\sqrt{d_{\max}-r_{\sigma}}+c_{4}\sigma_{\max}^{*}+c_{5}\sqrt{r_{\sigma}}\right).$ (115)
All the explicit expressions of the constants above can be found in Table 1.
Proof.
By the fact that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\epsilon_{\bm{\sigma}^{*}}$, together with 3, we obtain that there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for each $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (48) holds. Combining this with Section B.4 in Appendix D, we have that (86) and (70) (resp., (72)) hold when $L=2$ (resp., $L\geq 3$). Moreover, it follows from 4 and 5 that (73), (89), (94) hold, where $\tilde{\bm{U}}_{L}$, $\tilde{\bm{V}}_{1}$, and $\hat{\bm{T}}$ are respectively defined in (91), (92), and (93). Recall that $\bm{U}_{l}\in\mathcal{O}^{d_{l}}$ and $\bm{V}_{l}\in\mathcal{O}^{d_{l-1}}$ for all $l\in[L]$ are introduced in (3.2.1). For ease of exposition, we define $\bm{\Sigma}_{l}^{*}:=\mathrm{BlkDiag}(\mathrm{diag}(\bm{\sigma}^{*}),\bm{0})\in\mathbb{R}^{d_{l}\times d_{l-1}}$ and
$\hat{\bm{W}}_{1}:=\bm{U}_{1}\bm{\Sigma}_{1}^{*}\tilde{\bm{V}}_{1}^{T},\qquad\hat{\bm{W}}_{l}:=\bm{U}_{l}\bm{\Sigma}_{l}^{*}\bm{V}_{l}^{T},\ l=2,\dots,L-1,\qquad\hat{\bm{W}}_{L}:=\tilde{\bm{U}}_{L}\bm{\Sigma}_{L}^{*}\bm{V}_{L}^{T}.$ (116)
Now, we compute
$\|\hat{\bm{W}}_{1}-\bm{W}_{1}\|_{F}$ $=\left\|\bm{U}_{1}\bm{\Sigma}_{1}^{*}\tilde{\bm{V}}_{1}^{T}-\bm{U}_{1}\bm{\Sigma}_{1}\bm{V}_{1}^{T}\right\|_{F}$
$\leq\left\|\bm{U}_{1}\bm{\Sigma}_{1}^{*}\tilde{\bm{V}}_{1}^{T}-\bm{U}_{1}\bm{\Sigma}_{1}^{*}\bm{V}_{1}^{T}\right\|_{F}+\left\|\bm{U}_{1}\bm{\Sigma}_{1}^{*}\bm{V}_{1}^{T}-\bm{U}_{1}\bm{\Sigma}_{1}\bm{V}_{1}^{T}\right\|_{F}$
$\leq\sigma_{\max}^{*}\|\tilde{\bm{V}}_{1}-\bm{V}_{1}\|_{F}+\|\bm{\Sigma}_{1}^{*}-\bm{\Sigma}_{1}\|_{F}$
$\leq\left(c_{4}\sigma_{\max}^{*}+c_{3}\sqrt{d_{\max}-r_{\sigma}}+c_{5}\sqrt{r_{\sigma}}\right)\|\nabla G(\bm{W})\|_{F},$ (117)
where the last inequality uses (89), (73), and (94). Using the same argument, we obtain
$\|\hat{\bm{W}}_{L}-\bm{W}_{L}\|_{F}\leq\left(c_{4}\sigma_{\max}^{*}+c_{3}\sqrt{d_{\max}-r_{\sigma}}+c_{5}\sqrt{r_{\sigma}}\right)\|\nabla G(\bm{W})\|_{F}.$ (118)
For all $l=2,\dots,L-1$, we compute
[…]
This, together with (3.2.2) and (118), yields
$\|\bm{W}-\hat{\bm{W}}\|_{F}\leq\sqrt{L}\left(c_{4}\sigma_{\max}^{*}+c_{3}\sqrt{d_{\max}-r_{\sigma}}+c_{5}\sqrt{r_{\sigma}}\right)\|\nabla G(\bm{W})\|_{F}.$ (119)
Next, we further define
$\bm{T}_{l}:=\mathrm{BlkD}(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)}),\ l=2,\dots,L,\ \bm{Q}_{L}:=\bm{V}_{L},$ (120)
$\bm{W}_{1}^{*}:=\bm{Q}_{2}\bm{\Sigma}_{1}^{*}\hat{\bm{T}}^{T}\tilde{\bm{V}}_{1}^{T},\ \bm{W}_{l}^{*}:=\bm{Q}_{l+1}\bm{\Sigma}_{l}^{*}\bm{Q}_{l}^{T},\ l=2,\dots,L-1,\ \bm{W}_{L}^{*}:=\tilde{\bm{U}}_{L}\bm{\Sigma}_{L}^{*}\bm{Q}_{L}^{T},$ (121)
$\textstyle\bm{Q}_{l}:=\bm{V}_{l}\mathrm{BlkD}\left(\left(\prod_{k=l+1}^{L}\bm{T}_{k}^{(1)}\right),\dots,\left(\prod_{k=l+1}^{L}\bm{T}_{k}^{(p)}\right),\bm{I}_{d_{l-1}-r_{\sigma}}\right),\ l=2,\dots,L-1.$ (122)
Using 1, one can verify that $\bm{W}^{*}=(\bm{W}_{1}^{*},\bm{W}_{2}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}.$ For each $l=2,\dots,L-2$, we compute
$\|\hat{\bm{W}}_{l}-\bm{W}_{l}^{*}\|_{F}$ $=\left\|\bm{U}_{l}\bm{\Sigma}_{l}^{*}\bm{V}_{l}^{T}-\bm{Q}_{l+1}\bm{\Sigma}_{l}^{*}\bm{Q}_{l}^{T}\right\|_{F}$
[…]
where the last equality uses the diagonal block forms of $\bm{T}_{l}$ and $\bm{\Sigma}_{l}^{*}$. Using (116) and (121), we compute
$\|\hat{\bm{W}}_{1}-\bm{W}_{1}^{*}\|_{F}$ $=\|\bm{U}_{1}\bm{\Sigma}_{1}^{*}\tilde{\bm{V}}_{1}^{T}-\bm{Q}_{2}\bm{\Sigma}_{1}^{*}\hat{\bm{T}}^{T}\tilde{\bm{V}}_{1}^{T}\|_{F}=\|\bm{V}_{2}^{T}\bm{U}_{1}\bm{\Sigma}_{1}^{*}-\bm{V}_{2}^{T}\bm{Q}_{2}\bm{\Sigma}_{1}^{*}\hat{\bm{T}}^{T}\|_{F}$
$\leq\|\bm{V}_{2}^{T}\bm{U}_{1}\bm{\Sigma}_{1}^{*}-\bm{T}_{2}^{T}\bm{\Sigma}_{1}^{*}\|_{F}+\|\bm{T}_{2}^{T}\bm{\Sigma}_{1}^{*}-\bm{V}_{2}^{T}\bm{Q}_{2}\bm{\Sigma}_{1}^{*}\hat{\bm{T}}^{T}\|_{F}$
$=\|(\bm{U}_{1}^{T}\bm{V}_{2}-\bm{T}_{2})^{T}\bm{\Sigma}_{1}^{*}\|_{F}+\|\bm{\Sigma}_{1}^{*}-\bm{T}_{2}\bm{V}_{2}^{T}\bm{Q}_{2}\bm{\Sigma}_{1}^{*}\hat{\bm{T}}^{T}\|_{F}$
[…]
where the last equality follows from the froms of $\bm{T}_{l}$ in (120), $\bm{Q}_{2}$ in (122), and $\hat{\bm{T}}$ in (93). This, together with (3.2.2) and the fact that $\bm{W}_{L}^{*}=\hat{\bm{W}}_{L}$ due to $\bm{Q}_{L}=\bm{V}_{L}$, yields
$\|\hat{\bm{W}}-\bm{W}^{*}\|_{F}\leq\frac{9\sigma_{\max}^{*2}\sqrt{L-1}}{4\delta_{\sigma}\lambda\sigma^{*}_{\min}}\|\nabla G(\bm{W})\|_{F}.$ (124)
Since $\bm{W}^{*}\in\mathcal{W}_{\bm{\sigma}^{*}}$, we have $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\|\bm{W}-\bm{W}^{*}\|_{F}\leq\|\bm{W}-\hat{\bm{W}}\|_{F}+\|\hat{\bm{W}}-\bm{W}^{*}\|_{F}.$ This, together with (119) and (124), implies (113). ∎
Using 3 and the connection of Problems (1) and (12), we are now ready to prove 2.
Proof.
Proof of 2. According to 3, we conclude that for any $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$, (113) holds for $\mathcal{W}_{\bm{\sigma}^{*}}$, where the constants $\epsilon_{\bm{\sigma}^{*}}$ and $\kappa_{\bm{\sigma}^{*}}$ (see (114) and (115)) depend on $\bm{\sigma}^{*}$. Recalling the definition of $\mathcal{A}$ in (26), and noting that a polynomial has only finitely many roots, it follows that $\mathcal{A}_{\rm sort}$ defined in (29) is finite. Consequently, by (29), we can derive a corresponding result for $\mathcal{W}_{G}$ that does not depend on $\bm{\sigma}^{*}$. In addition, we take the analysis about $\bm{\sigma}^{*}=\bm{0}$ in 2 into consideration. Specifically, by letting
$\kappa:=\max_{\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}}\kappa_{\bm{\sigma}^{*}}\quad\text{and}\quad\epsilon:=\min_{\bm{\sigma}*\in\mathcal{A}_{\rm sort}}\epsilon_{\bm{\sigma}^{*}},$ (125)
we have $\mathrm{dist}(\bm{W},\mathcal{W}_{G})\leq\kappa\|\nabla G(\bm{W})\|_{F}$ for all $\mathrm{dist}(\bm{W},\mathcal{W}_{G})\leq\epsilon.$ Using this and Lemma 1(ii), (9) holds with
$\kappa_{1}={\kappa\lambda}/{\lambda_{\min}}\quad\text{and}\quad\epsilon_{1}={\epsilon}/{\sqrt{\lambda_{\max}}}.$ (126)
4 Experimental Results
In this section, we conduct experiments under different settings to validate our theoretical results. Specifically, we employ GD to train deep networks using the PyTorch library. We terminate the algorithm when the squared gradient norm satisfies $\|\nabla F(\bm{W}^{k})\|_{F}^{2}\leq 10^{-6}$ and the function value change satisfies $|F(\bm{W}^{k})-F(\bm{W}^{k-1})|\leq 10^{-7}$ for all $k=1,2,\dots$, where $\bm{W}^{k}$ is the $k$-th iterate.
4.1 Linear Convergence to Critical Points
In this subsection, we investigate the convergence behavior of GD to different critical points of Problem (1). In our experiments, we set $d_{L}=20$, $d_{0}=10$, and $d_{l}=32$ for each $l=2,\dots,L-1$. Then, we i.i.d. sample each entry of $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$ from the standard Gaussian distribution, i.e., $y_{ij}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. Moreover, we set the regularization parameters $\lambda_{l}=10^{-4}$ for all $l\in[L]$ and the learning rate $4.5\times 10^{-4}$. To ensure convergence to different critical points, we initialize the weight matrices in the neighborhood of two different critical points of Problem (1). Specifically, we apply the SVD to $\bm{Y}$ and solve the equation (6) to get its roots. Using these results and 1, we can respectively construct an optimal solution, denoted by $\bm{W}_{\rm opt}^{*}$, and a non-optimal critical point, denoted by $\bm{W}_{\rm crit}^{*}$, of Problem (1). Then, we set $\bm{W}^{0}=\bm{W}^{*}+0.01\bm{\Delta}$, where $\bm{W}^{*}$ is $\bm{W}_{\rm opt}^{*}$ or $\bm{W}_{\rm crit}^{*}$ and each entry of $\bm{\Delta}$ is i.i.d. sampled from the standard Gaussian distribution. For each initialization, we run GD for solving Problem (1) with different depths $L\in\{2,4,6\}$.
(a) Training loss
(b) Squared gradient norm
(c) Training loss
(d) Squared gradient norm
Figure 1: Linear convergence of GD to a critical point. In (a) and (b), $\{\bm{W}^{k}\}$ converges to an optimal solution when initialized near an optimal solution; in (c) and (d), $\{\bm{W}^{k}\}$ converges to a non-optimal critical point for $L=4,6$, and to an optimal solution for $L=2$ when initialized near a non-optimal critical point. Here $F^{\rm end}$ denotes the training loss when the stopping criterion is met.
Initialization Objective values $L=2$ $L=4$ $L=6$
Initialization near global optimal solution $F(\bm{W}_{\rm opt}^{*})$ $9.2900\times 10^{-3}$ $8.1723\times 10^{-3}$ $9.5264\times 10^{-3}$
$F(\bm{W}^{\rm end})$ $9.2985\times 10^{-3}$ $8.2020\times 10^{-3}$ $9.5741\times 10^{-3}$
Initialization near non-optimal critical point $F(\bm{W}^{*}_{\rm cri})$ $6.8038\times 10^{-1}$ $6.7906\times 10^{-1}$ $6.8022\times 10^{-1}$
$F(\bm{W}^{\rm end})$ $9.2980\times 10^{-3}$ $6.7909\times 10^{-1}$ $6.8027\times 10^{-1}$
Table 2: The function values at the final iterate $\bm{W}^{\rm end}$ and the critical point $\bm{W}^{*}$.
We respectively plot the function value gap $F(\bm{W}^{k})-F(\bm{W}^{\rm end})$ and the squared gradient norm $\|\nabla F(\bm{W}^{k})\|_{F}^{2}$ against the iteration number in Figure 1(a) and (b) (resp., Figure 1(c) and (d)), where $\bm{W}^{\rm end}$ denotes the last iterate of the GD. We also report the function values of the final iterate, in different settings in Table 2. As observed from Figure 1 and Table 2, GD converges to an optimal solution at a linear rate for solving Problem (1) with different network depths, and similarly, it converges to a non-optimal critical point at a linear rate for $L=4,6$. We should point out that Problem (1) only has global optima and strict saddle points when $L=2$, as shown in [6]. Therefore, even if initialized in the vicinity of a non-optimal critical point, i.e., strict saddle point, the algorithm will ultimately converge to a global optimal point almost surely [27]. This aligns with our linear convergence analysis for all critical points of Problem (1) in 6 in Appendix C, which leverages the error bound of Problem (1), and thereby supports 2.
4.2 Convergence Behavior in General Settings
In this subsection, we investigate the convergence behavior of GD in more general setups extending beyond linear networks. Specifically, the network depth is fixed as $L=4$ and network widths at different layers are set as follows: $d_{L}=16$, $d_{0}=10$, and $d_{l}=32$ for each $l=2,\dots,L-1$. The regularization parameter is set as $\lambda_{l}=5\times 10^{-5}$ for all $l\in[L]$ and the learning rate of GD is set as $10^{-3}$. We use the default initialization scheme in PyTorch to initialize the weights for GD. The data matrix $\bm{X}$ is generated according to a uniform distribution using the xavier_uniform_ function in PyTorch, while the target matrix $\bm{Y}$ is generated using the same approach as described in Section 4.1. Below, we outline the different setups used in our experiments.
General data input. We first consider general data inputs $\bm{X}_{1},\bm{X}_{2},\bm{X}_{3}$ instead of orthogonal inputs $\bm{X}$. Then, we apply GD to optimize Problem (1) with a linear network $\bm{W}_{L}\ldots\bm{W}_{1}\bm{X}$ for each of these different data matrices.
Linear networks with bias. We next study the regularized loss of deep linear networks with bias terms:
$\min\limits_{\{\bm{W}_{l},\bm{b}_{l}\}}\;\sum_{i=1}^{N}\left\|\bm{W}_{L}(\bm{W}_{L-1}\cdots(\bm{W}_{1}\bm{x}_{i}+\bm{b}_{1})+\bm{b}_{L-1})+\bm{b}_{L}-\bm{y}_{i}\right\|^{2}+\sum_{l=1}^{L}\lambda_{l}\left(\left\|\bm{W}_{l}\right\|_{F}^{2}+\left\|\bm{b}_{l}\right\|^{2}\right),$
where $\bm{x}_{i}$ and $\bm{y}_{i}$ respectively denote the $i$-th column of $\bm{X}$ and $\bm{Y}$. We apply GD to solve the above problem when the input is either an identity matrix or general input data.
Deep nonlinear networks. Finally, we study the performance of GD for solving the regularized loss of deep nonlinear networks with different activation functions:
$\min\limits_{\{\bm{W}_{l},\bm{b}_{l}\}}\ \sum_{i=1}^{N}\left\|\bm{W}_{L}\sigma\left(\bm{W}_{L-1}\cdots\sigma\left(\bm{W}_{1}\bm{x}_{i}+\bm{b}_{1}\right)+\bm{b}_{L-1}\right)+\bm{b}_{L}-\bm{y}_{i}\right\|_{F}^{2}+\sum_{l=1}^{L}\lambda_{l}\left(\left\|\bm{W}_{l}\right\|_{F}^{2}+\left\|\bm{b}_{l}\right\|_{F}^{2}\right),$
where $\sigma(\cdot)$ denotes an activation function. We set $\bm{X}$ as the identity matrix and use GD for solving the above problem when the activation functions are chosen as ReLU, Leaky ReLU, and tanh, respectively.
(a) General data input
(b) Linear networks with bias
(c) Deep nonlinear networks
Figure 2: Linear convergence of GD under different settings: (a) Three different matrices $\bm{X}_{1}$, $\bm{X}_{2}$, and $\bm{X}_{3}$, with condition numbers of 43.43, 36.04, and 16.36, respectively. (b) Identity matrix, $\bm{X}_{1}$, and $\bm{X}_{2}$, with condition numbers of 1, 4.82, and 14.62, respectively. (c) Identity matrix as data input with ReLU, Leaky ReLU, and tanh as activation functions. $F^{\rm end}$ denotes the training loss at the final step when the stop condition is triggered.
For the above three different settings, we plot the function value gap $F(\bm{W}^{k})-F(\bm{W}^{\text{end}})$ against the iteration number in Figure 2. It is observed from Figure 2 that GD converges to a solution at a linear rate across these settings. This consistent behavior leads us to conjecture that the error-bound condition may hold for deep networks in more general scenarios. Additionally, we observe that the number of iterations required to meet the stopping criterion generally increases as the condition number of the input data becomes larger. Exploring this phenomenon presents an interesting direction for future research.
5 Conclusions
In this paper, we studied the regularized squared loss of deep linear networks and proved its error bound, a regularity condition that characterizes local geometry around the critical point set. This result is not only theoretically significant but also lays down the foundation for establishing strong convergence guarantees for various methods for solving Problem (1). To establish the error bound, we explicitly characterized the critical point set of (1) and developed new analytic techniques to show the error bound, which may be of independent interest. Our numerical results across different settings provide strong support for our theoretical findings. One future direction is to extend our analysis to deep linear networks with more general data input $\bm{X}$ and loss functions. Another interesting direction is to investigate the regularized loss of deep nonlinear networks.
Appendix A Showing the Necessity of Assumption 2
We claim that the error bound of $F$ (see Problem (1)) holds if and only if the error bound of $G$ (see Problem (12)) holds. Indeed, it follows from (iii) in Lemma 1 that the “if” direction holds. Now, it remains to show the “only if” direction. Suppose that the error bound of Problem (1) holds (see (9)). Using the proof setup in Lemma 1, we have
[…]
Therefore, we prove the claim. Based on the error-bound equivalence between $F$ and $G$, establishing the necessity of Assumption 2 for $F$ is equivalent to doing so for $G$.
A.1 The Case $L=2$
For ease of exposition, we denote $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}:=\mathcal{W}_{\rm sort(\bm{\sigma}^{*})}$ for each $\bm{\sigma}^{*}\in\mathcal{A}$, where $\mathrm{sort}(\cdot)$ is a sorting function that arranges the elements of a vector in decreasing order.
Lemma A.1.
Suppose that $L=2$ and (7) does not hold. For any $\bm{\sigma}^{*}\in\cal A$, the error bound for the critical point set $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}$ of Problem (12) fails to hold.
Proof.
Since (7) does not hold, there exists $i\in[r_{Y}]$ such that $y_{i}=\sqrt{\lambda}$. Now, we define $\bm{W}(t):=\left(\bm{W}_{1}(t),\bm{W}_{2}(t)\right)$ as follows:
$\textstyle\begin{cases}&\bm{W}_{1}(t)=\bm{Q}_{2}\bm{\Sigma}_{1}(t)\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ &\bm{W}_{2}(t)=\mathrm{BlkD}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{2}(t)\bm{Q}_{2}^{T},\\ &\bm{\Sigma}_{l}(t)=\mathrm{BlkD}\left(\mathrm{diag}(\bm{\sigma}^{*}+t\bm{e}_{i}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\forall l=1,2,\\ &\bm{O}_{i}\in\mathcal{O}^{h_{i}},\forall i\in[p_{Y}],\bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{2}-r_{Y}}.\end{cases}$
According to 1, one can verify that $\bm{W}(0)\in\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}$. Using $y_{i}=\sqrt{\lambda}$, we obtain
$\sigma^{*}_{i}\in\{x:x^{3}-\sqrt{\lambda}y_{i}x+\lambda x=0,\ x\geq 0\}=\{0\}.$
Then, we compute
[…]
where the second equality follows from $y_{i}=\sqrt{\lambda}$. Moreover, we have
$\mathrm{dist}^{2}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})$ $=\min_{(\bm{W}_{1},\bm{W}_{2})\in\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}}\|\bm{W}_{1}(t)-\bm{W}_{1}\|_{F}^{2}+\|\bm{W}_{L}(t)-\bm{W}_{2}\|_{F}^{2}$
$\geq\|\bm{\Sigma}_{1}(t)-\bm{\Sigma}_{1}(0)\|_{F}^{2}+\|\bm{\Sigma}_{2}(t)-\bm{\Sigma}_{2}(0)\|_{F}^{2}=2t^{2},$ (128)
where the first inequality uses Mirsky’s inequality (see Lemma D.8). When $t\leq{\delta_{\sigma}}/{3}$, we have $\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})\leq\|\bm{W}(t)-\bm{W}(0)\|_{F}\leq{\sqrt{2}\delta_{\sigma}}/{3}.$ Using this and 1(ii), we have for each $\bar{\bm{\sigma}}^{*}\in\mathcal{A}$ satisfying $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}\neq\widehat{\mathcal{W}}_{\bar{\bm{\sigma}}^{*}}$,
$\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bar{\bm{\sigma}}^{*}})\geq\mathrm{dist}(\widehat{\mathcal{W}}_{\bm{\sigma}^{*}},\widehat{\mathcal{W}}_{\bar{\bm{\sigma}}^{*}})-\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})\geq\left(1-\frac{\sqrt{2}}{3}\right)\delta_{\sigma}>\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}).$
Consequently, we obtain $\mathrm{dist}(\bm{W}(t),\mathcal{W}_{G})=\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})$. This, together with (127) and (128), implies $\|\nabla G(\bm{W}(t))\|_{F}=2\sqrt{2}|t|^{3}\leq 2^{3/4}\mathrm{dist}^{\frac{3}{2}}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})=2^{3/4}\mathrm{dist}^{\frac{3}{2}}(\bm{W}(t),\mathcal{W}_{G}).$ Obviously, the error bound does not hold when $t\to 0$. ∎
A.2 The Case $L\geq 3$
We first present the following lemma to derive the equivalent condition to (8).
Lemma A.2.
When $L\geq 3$, (8) holds if and only if $\varphi^{\prime}(\sigma^{*})\neq 0$ for all $\sigma^{*}\in\mathcal{Y}\setminus\{0\}$, where $\cal Y$ and $\varphi(\cdot)$ are defined in (28) and (85), respectively. When $L=2$, we have $\varphi^{\prime}(\sigma^{*})\neq 0$.
Proof.
First, consider the case where $L\geq 3$. Suppose that there exists $\sigma^{*}\in\mathcal{Y}\setminus\{0\}$ such that $\varphi^{\prime}(\sigma^{*})=0$. Note that $\varphi^{\prime}(\sigma^{*})=0$ is equivalent to $\varphi^{\prime}(\sigma^{*})=L\sigma^{*(L-1)}/\sqrt{\lambda}+\sqrt{\lambda}(2-L)\sigma^{*(1-L)}=0.$ Using $\sigma^{*}\in\mathcal{Y}$, there exists $j\in[d_{\min}]$ such that $\sigma^{*(2L-1)}+\lambda\sigma^{*}=\sqrt{\lambda}\sigma^{*(L-1)}y_{j}$. Combining the above two equations, we obtain
$y_{j}=\left(\left(\frac{L-2}{L}\right)^{\frac{L}{2(L-1)}}+\left(\frac{L}{L-2}\right)^{\frac{L-2}{2(L-1)}}\right)\lambda^{\frac{1}{2(L-1)}}.$ (129)
This implies that (8) does not hold.
Conversely, suppose that (8) does not hold. This implies that there exists $j\in[d_{\min}]$ such that (129) holds. Then, one can verify that $x^{2L-1}-\sqrt{\lambda}y_{j}x^{L-1}+\lambda x=0$ has a positive root $x^{*}=\left({\lambda(L-1)}/{L}\right)^{\frac{1}{2(L-1)}}$ and $\varphi^{\prime}(x^{*})=0$. Using the definition of $\mathcal{Y}$, we have $x^{*}\in\mathcal{Y}$.
When $L=2$, we compute $\varphi^{\prime}(\sigma^{*})=2\sigma^{*}/\sqrt{\lambda}>0$ for all $\sigma^{*}\in\mathcal{Y}\setminus\{0\}.$ So, $\varphi^{\prime}(\sigma^{*})\neq 0$ holds trivially. ∎
Lemma A.3.
Suppose that $L\geq 3$ and (8) does not hold. There exists $\bm{\sigma}^{*}\in\cal A$ such that the error bound for the critical point set $\widehat{\mathcal{W}}_{\bm{\sigma}^{*}}$ of Problem (12) fails to hold.
Proof.
According to Lemma A.2 and the fact that (8) does not hold, there exists $\bm{\sigma}^{*}\in\cal A$ such that $\varphi^{\prime}(\sigma_{i}^{*})=0$ for some $i\in[d_{\min}]$. This implies $f(\sigma^{*}_{i}):=(\sigma^{*}_{i})^{2L-1}-\sqrt{\lambda}y_{i}(\sigma^{*}_{i})^{L-1}+\lambda\sigma^{*}_{i}=0$ and $\varphi(\sigma^{*}_{i})=y_{i}$. Note that $f(x)=\sqrt{\lambda}x^{L-1}\varphi(x)-\sqrt{\lambda}y_{i}x^{L-1}$ and we compute $f^{\prime}(\sigma_{i}^{*})=\sqrt{\lambda}(\sigma_{i}^{*})^{L-1}\varphi^{\prime}(\sigma_{i}^{*})=0$. Now, we define $\bm{W}(t)=(\bm{W}_{1}(t),\cdots,\bm{W}_{L}(t))$ as follow:
$\begin{cases}&\bm{W}_{1}(t)=\bm{Q}_{2}\bm{\Sigma}_{1}(t)\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right),\\ &\bm{W}_{l}(t)=\bm{Q}_{l+1}\bm{\Sigma}_{l}(t)\bm{Q}_{l}^{T},\ l=2,\dots,L-1,\\ &\bm{W}_{L}(t)=\mathrm{BlkD}\left(\bm{O}_{1}^{T},\dots,\bm{O}_{p_{Y}}^{T},\widehat{\bm{O}}_{p_{Y}+1}^{T}\right)\bm{\Sigma}_{L}(t)\bm{Q}_{L}^{T},\\ &\bm{\Sigma}_{l}(t)=\mathrm{BlkD}\left(\mathrm{diag}(\bm{\sigma}^{*}+t\bm{e}_{i}),\bm{0}\right)\in\mathbb{R}^{d_{l}\times d_{l-1}},\ \forall l\in[L],\\ &\bm{O}_{i}\in\mathcal{O}^{h_{i}},\forall i\in[p_{Y}],\ \bm{O}_{p_{Y}+1}\in\mathcal{O}^{d_{0}-r_{Y}},\widehat{\bm{O}}_{p_{Y}+1}\in\mathcal{O}^{d_{L}-r_{Y}},\ \bm{Q}_{l}\in\mathcal{O}^{d_{l-1}},\ l=2,\dots,L.\end{cases}$
It follows from 1 that $\bm{W}(0)\in\mathcal{W}_{G}$. Therefore, we obtain for all $l\in[L]$,
[…]
Applying the Taylor expansion to $f(\sigma^{*}_{i}+t)$ at $\sigma^{*}$, together with $f(\sigma^{*}_{i})=0$ and $f^{\prime}(\sigma^{*}_{i})=0$, yields that when $t\to 0$, $\|\nabla G(\bm{W}(t))\|_{F}=O(t^{2}).$ We also note that $\mathrm{dist}^{2}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})=\|\bm{W}_{1}(t)-\bm{W}_{1}^{*}\|_{F}^{2}+\cdots+\|\bm{W}_{L}(t)-\bm{W}_{L}^{*}\|_{F}^{2}\geq Lt^{2},$ where the inequality follows from Weyl’s inequality. Using the same argument in Lemma A.1, we conclude that $\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})=\mathrm{dist}(\bm{W}(t),\mathcal{W}_{G})$ when $t$ is sufficiently small. Then we have $\left\|\nabla G\left(\bm{W}(t)\right)\right\|_{F}=O\left(\mathrm{dist}^{2}(\bm{W}(t),\mathcal{W}_{G})\right)$, which implies that the error bound fails to hold. ∎
Remark 1.
Under Assumption 1, it follows from 2 that Assumption 2 is a sufficient condition for the error bound to hold for the critical point set $\mathcal{W}_{G}$. According to Lemma A.1 and Lemma A.3, we conclude that Assumption 2 is also a necessary condition. Therefore, we establish the necessary and sufficient condition unde which the error bound holds.
Appendix B Supplementary Results and Proofs for Section 3
B.1 Proof of 1
Proof.
(i) Obviously, the “only if” direction is trivial since $\bm{\sigma}\in\mathcal{W}_{\bm{\sigma},\bm{\Pi}}$ and $\bm{\sigma}^{\prime}\in\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}$ share the same non-increasing singular values. It remains to prove the “if” direction. Suppose that $(\bm{\sigma},\bm{\Pi})\in\mathcal{B}$ and $(\bm{\sigma}^{\prime},\bm{\Pi}^{\prime})\in\mathcal{B}$ satisfy $\bm{\sigma}=\bm{\sigma}^{\prime}$. There exist $\bm{a},\bm{a}^{\prime}\in\mathcal{A}$ such that
$\bm{a}=\bm{\Pi}^{T}\bm{\sigma},\ \bm{a}^{\prime}=\bm{\Pi}^{\prime T}\bm{\sigma}^{\prime}.$ (130)
This, together with $\bm{\sigma}=\bm{\sigma}^{\prime}$, implies that $\bm{a},\bm{a}^{\prime}$ have the same positive elements but in a different order. Consider the following equation from $\mathcal{A}$: $x_{i}^{2L-1}-\sqrt{\lambda}y_{i}x_{i}^{L-1}+\lambda x_{i}=0,\ x_{i}\geq 0,\ \forall i\in[d_{\min}].$ Note that if $x_{i}=x_{j}>0$, we have $y_{i}=y_{j}$, which implies that if $y_{i}\neq y_{j}$, we have $x_{i}\neq x_{j}$ when $x_{i},x_{j}>0$. Since $\bm{\sigma}=\bm{\sigma}^{\prime}$ and each element of $\bm{a}$ and $\bm{a}^{\prime}$ is obtained by solving the above equation, along with the fact that different $y_{i}$’s correspond to equations with no common positive root and the partition of $(y_{1},\ldots,y_{d_{\min}})$ in (4), yields that the elements of $\bm{a}$ and $\bm{a}^{\prime}$ in each partition of the form $\bm{Y}$ differ only in their order. Therefore,there exist $\bm{P}_{i}\in\mathcal{P}^{h_{i}}$ for all $i\in[p_{Y}]$ and $\bm{P}_{p_{Y}+1}\in\mathcal{P}^{d_{\min}-r_{Y}}$ such that $\mathrm{diag}(\bm{a})=\mathrm{BlkDiag}(\bm{P}_{1},\ldots,\bm{P}_{p_{Y}},\bm{P}_{p_{Y}+1})\mathrm{diag}(\bm{a}^{\prime})\mathrm{BlkDiag}(\bm{P}_{1}^{T},\ldots,\bm{P}_{p_{Y}}^{T},\bm{P}_{p_{Y}+1}^{T}).$ Substituting (130) into the above equality, together with $\bm{\sigma}=\bm{\sigma}^{\prime}$, yields
$\bm{\Pi}^{T}\mathrm{diag}(\bm{\sigma})\bm{\Pi}=\mathrm{BlkDiag}(\bm{P}_{1},\ldots,\bm{P}_{p_{Y}},\bm{P}_{p_{Y}+1})\bm{\Pi}^{\prime T}\mathrm{diag}(\bm{\sigma})\bm{\Pi}^{\prime}\mathrm{BlkDiag}(\bm{P}_{1}^{T},\ldots,\bm{P}_{p_{Y}}^{T},\bm{P}_{p_{Y}+1}^{T}).$ (131)
Let $(\bm{W}_{1},\cdots,\bm{W}_{L})\in\mathcal{W}_{\bm{\sigma},\bm{\Pi}}$ be arbitrary. For ease of exposition, let
$\tilde{\bm{Q}}_{l}:=\bm{Q}_{l}\mathrm{BlkDiag}(\bm{\Pi},\bm{I})\mathrm{BlkDiag}\left(\bm{P}_{1},\ldots,\bm{P}_{p_{Y}},\bm{P}_{p_{Y}+1},\bm{I}\right)\mathrm{BlkDiag}(\bm{\Pi}^{{}^{\prime}T},\bm{I})\in\mathcal{O}^{d_{l-1}},\ \forall l\in\{2,\dots,L\}.$ (132)
Moreover, we have $\bm{\Sigma}_{l}\mathrm{BlkDiag}(\bm{\Pi},\bm{I})=\mathrm{BlkDiag}(\mathrm{diag}(\bm{\sigma})\bm{\Pi},\bm{0}),\ \forall l\in[L].$ Then, we have
$\textstyle\bm{W}_{1}$ $=\bm{Q}_{2}\bm{\Sigma}_{1}\mathrm{BlkD}\left(\bm{\Pi},\bm{I}\right)\mathrm{BlkD}\left(\bm{O}_{1},\dots,\bm{O}_{p_{Y}},\bm{O}_{p_{Y}+1}\right)$
$=\tilde{\bm{Q}}_{2}\bm{\Sigma}_{1}\mathrm{BlkD}\left(\bm{\Pi}^{\prime},\bm{I}\right)\mathrm{BlkDiag}(\bm{P}_{1}^{T}\bm{O}_{1},\ldots,\bm{P}_{p_{Y}}^{T}\bm{O}_{p_{Y}},\mathrm{BlkDiag}(\bm{P}_{p_{Y}+1}^{T},\bm{I}_{d_{0}-d_{\min}})\bm{O}_{p_{Y}+1}),$
where the second equality uses (131) and (132). Using the same argument, we have
$\bm{W}_{L}=\mathrm{BlkD}\left(\bm{O}_{1}^{T}\bm{P}_{1},\dots,\bm{O}_{p_{Y}}^{T}\bm{P}_{Y},\widehat{\bm{O}}_{p_{Y}+1}^{T}\mathrm{BlkDiag}(\bm{P}_{p_{Y}+1},\bm{I})\right)\mathrm{BlkD}\left(\bm{\Pi}^{\prime T},\bm{I}\right)\bm{\Sigma}_{L}\tilde{\bm{Q}}_{L}^{T},$
$\bm{W}_{l}=\tilde{\bm{Q}}_{l+1}\bm{\Sigma}_{l}\tilde{\bm{Q}}_{l}^{T},\ l=2,\dots,L-1.$
Therefore, we obtain that $(\bm{W}_{1},\ldots,\bm{W}_{L})\in\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}$ and thus $\mathcal{W}_{\bm{\sigma},\bm{\Pi}}\subseteq\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}$. Applying the same argument, we also have $\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}\subseteq\mathcal{W}_{\bm{\sigma},\bm{\Pi}}$. Therefore, we have $\mathcal{W}_{\bm{\sigma},\bm{\Pi}}=\mathcal{W}_{\bm{\sigma}^{\prime},\bm{\Pi}^{\prime}}$.
[…]
B.2 Auxiliary Lemma for Proving Lemma 3
Lemma B.4.
Let $a>0$ be a constant, $\bm{\Sigma}\in\mathbb{R}^{n\times n}$ be a diagonal matrix, and $\bm{Q}\in\mathcal{O}^{n}$ be an orthogonal matrix. Then, if $\|\bm{\Sigma}-a\bm{I}\|\leq a/2$, we have $\|\bm{Q}\bm{\Sigma}^{2}-\bm{\Sigma}^{2}\bm{Q}\|_{F}\geq a\|\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}\|_{F}.$
Proof.
For ease of exposition, let $\bm{\Delta}:=\bm{\Sigma}-a\bm{I}$. We compute
$\textstyle\|\bm{Q}\bm{\Sigma}^{2}-\bm{\Sigma}^{2}\bm{Q}\|_{F}$ $=\|\bm{Q}\left(\bm{\Delta}+a\bm{I}\right)^{2}-\left(\bm{\Delta}+a\bm{I}\right)^{2}\bm{Q}\|_{F}$
$=\left\|\bm{Q}\bm{\Delta}^{2}-\bm{\Delta}^{2}\bm{Q}+2a\left(\bm{Q}\bm{\Delta}-\bm{\Delta}\bm{Q}\right)\right\|_{F}$
$\geq 2a\left\|\bm{Q}\bm{\Delta}-\bm{\Delta}\bm{Q}\right\|_{F}-\left\|\bm{Q}\bm{\Delta}^{2}-\bm{\Delta}^{2}\bm{Q}\right\|_{F}$
$\geq 2\left(a-\|\bm{\Delta}\|\right)\left\|\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}\right\|_{F}\geq a\left\|\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}\right\|_{F},$
where the second inequality follows from $\bm{Q}\bm{\Delta}-\bm{\Delta}\bm{Q}=\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}$ and $\|\bm{Q}\bm{\Delta}^{2}-\bm{\Delta}^{2}\bm{Q}\|_{F}=\|\bm{\Delta}(\bm{\Delta}\bm{Q}-\bm{Q}\bm{\Delta})+(\bm{\Delta}\bm{Q}-\bm{Q}\bm{\Delta})\bm{\Delta}\|_{F}\leq 2\|\bm{\Delta}\|\|\bm{Q}\bm{\Sigma}-\bm{\Sigma}\bm{Q}\|_{F}$, and the last inequality uses $\|\bm{\Delta}\|\leq a/2$. ∎
B.3 Auxiliary Lemma and Proof for Proving 5
Lemma B.5.
Let $\bm{\sigma}^{*}\in\mathcal{A}_{\mathrm{sort}}\setminus\{\bm{0}\}$ and $i\in[r_{\sigma}]$ be arbitrary. Suppose that $|x-\sigma_{i}^{*}|\leq\min\left\{\delta_{1},\delta_{2},\frac{\delta_{\sigma}}{3}\right\},$ where $\delta_{1}$ and $\delta_{2}$ are defined in (87). Then we have
$\left|\varphi(x)-\varphi(\sigma_{i}^{*})\right|\leq\frac{\delta_{y}}{3}\text{ and }\ |\varphi^{\prime}(x)-\varphi^{\prime}(\sigma^{*}_{i})|\leq\frac{|\varphi^{\prime}(\sigma_{i}^{*})|}{2}.$ (133)
Proof.
According to the definition of $\delta_{\sigma}$ in (28), if $|x-\sigma_{i}^{*}|\leq\frac{\delta_{\sigma}}{3}$, then we have $\frac{2\sigma_{\min}^{*}}{3}\leq|x|\leq\frac{4\sigma_{\max}^{*}}{3}.$ Consequently, we obtain that
$|\varphi^{\prime}(x)|$ $=\Bigl|\frac{L}{\sqrt{\lambda}}x^{L-1}+\sqrt{\lambda}(2-L)x^{1-L}\Bigr|\leq\frac{L}{\sqrt{\lambda}}\left(\frac{4\sigma_{\max}^{*}}{3}\right)^{L-1}+(L-2)\sqrt{\lambda}\left(\frac{2\sigma_{\min}^{*}}{3}\right)^{1-L},$
$|\varphi^{\prime\prime}(x)|$ $=\Bigl|\frac{L(L-1)}{\sqrt{\lambda}}x^{L-1}+\sqrt{\lambda}(2-L)(1-L)x^{-L}\Bigr|$
$\leq\frac{L(L-1)}{\sqrt{\lambda}}\left(\frac{4\sigma_{\max}^{*}}{3}\right)^{L-1}+\sqrt{\lambda}(2-L)(1-L)\left(\frac{2\sigma_{\min}^{*}}{3}\right)^{-L}.$
Finally, by the mean value theorem, for all $x$ satisfying $|x-\sigma_{i}^{*}|\leq\min\{\delta_{1},\delta_{2},\frac{\delta_{\sigma}}{3}\}$, we obtain (133). ∎
B.3.1 Proof of (97) and (98)
Proof.
Using the form of $\bm{\Sigma}_{l}$ in (3.2.1) and (71) for $L=2$ (resp., (73) for $L\geq 3$), we have for each $l\in[L]$,
$\left\|\hat{\bm{\Sigma}}_{l}-\bm{\Sigma}_{l}\right\|_{F}\!=\!\!\left\|\bm{\Sigma}_{l}^{(p+1)}\right\|_{F}\!\!\leq\!c_{3}\sqrt{\min\{d_{l},d_{l-1}\}}\|\nabla G(\bm{W})\|_{F}\leq c_{3}\sqrt{d_{\max}}\|\nabla G(\bm{W})\|_{F},$ (134)
$\left\|(\hat{\bm{\Sigma}}_{l}\hat{\bm{\Sigma}}_{l}^{T})^{L-1}\hat{\bm{\Sigma}}_{l}-(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}\right\|_{F}=\left\|\left(\bm{\Sigma}_{l}^{(p+1)}\bm{\Sigma}_{l}^{(p+1)^{T}}\right)^{L-1}\bm{\Sigma}_{l}^{(p+1)}\right\|_{F}$
[…]
We first bound
$\left\|(\hat{\bm{\Sigma}}_{L}\hat{\bm{\Sigma}}_{L}^{T})^{L-1}\hat{\bm{\Sigma}}_{L}+\lambda\hat{\bm{\Sigma}}_{L}-\sqrt{\lambda}\bm{\Psi}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{0}\right)\right\|_{F}$
$\leq$ $\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{\Psi}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}+\lambda\|\hat{\bm{\Sigma}}_{L}-\bm{\Sigma}_{L}\|_{F}$
$+\left\|(\hat{\bm{\Sigma}}_{L}\hat{\bm{\Sigma}}_{L}^{T})^{L-1}\hat{\bm{\Sigma}}_{L}-(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}\right\|_{F}+\sqrt{\lambda}\|\bm{\Psi}\|\|\bm{B}_{p+1}\|_{F}$
$\leq$ $\left(c_{2}+\left(\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(L-1)}+\lambda\right)c_{3}\sqrt{d_{\max}}\right)\left\|\nabla G(\bm{W})\right\|_{F}+\sqrt{\lambda}y_{1}\|\bm{B}_{p+1}\|_{F}$ (136)
$\leq$ $\eta_{3}\left\|\nabla G(\bm{W})\right\|_{F},$ (137)
where the second inequality follows from (55), (59), (134), and (135), and the last uses (35) and (55). Using (55) and (140), we have
$\sum_{i=1}^{r_{\sigma}}\left\|\left(\sigma_{i}^{2L-1}(\bm{W}_{L})+\lambda\sigma_{i}(\bm{W}_{L})\right)\bm{e}_{i}-\sqrt{\lambda}\sigma_{i}^{L-1}(\bm{W}_{L})\bm{\Psi}\hat{\bm{T}}\bm{e}_{i}\right\|^{2}$
$\leq$ $\left\|(\hat{\bm{\Sigma}}_{L}\hat{\bm{\Sigma}}_{L}^{T})^{L-1}\hat{\bm{\Sigma}}_{L}+\lambda\hat{\bm{\Sigma}}_{L}-\sqrt{\lambda}\bm{\Psi}\hat{\bm{T}}\mathrm{BlkD}\left((\bm{\Sigma}_{L}^{(1)})^{L-1},\dots,(\bm{\Sigma}_{L}^{(p)})^{L-1},\bm{0}\right)\right\|_{F}^{2}$
$=$ $\left\|(\hat{\bm{\Sigma}}_{L}\hat{\bm{\Sigma}}_{L}^{T})^{L-1}\hat{\bm{\Sigma}}_{L}+\lambda\hat{\bm{\Sigma}}_{L}-\sqrt{\lambda}\bm{\Psi}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{0}\right)\right\|_{F}^{2}$
Dividing the above inequality by $\sqrt{\lambda}\,\sigma_{r_{\sigma}}^{L-1}(\bm{W}_{L})$ and using (137) and the auxiliary function $\varphi$ in (85) yields
[…]
which is just the desired inequality (97). By applying the same argument used to derive (137) using (58), we have
$\left\|(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T})^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{0}\right)\bm{\Psi}\right\|_{F}\leq\eta_{3}\left\|\nabla G(\bm{W})\right\|_{F}.$ (139)
Now, we claim
$\left\|\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\,\mathrm{BlkD}\left((\bm{\Sigma}_{1}^{(1)})^{L-1},\dots,(\bm{\Sigma}_{1}^{(p)})^{L-1},\bm{0}\right)\bm{\Psi}\hat{\bm{T}}\right\|_{F}\leq\eta_{4}\left\|\nabla G(\bm{W})\right\|_{F}.$ (140)
Applying a similar argument as in (138) to the rows of the above inequality yields (98).
In the rest of this proof, we show that (140) holds. For each $i\in[p]$, we define $\hat{\bm{T}}_{i}:=\prod_{l=2}^{L}\bm{T}_{l}^{(i)}$. For each $i\in[p]$, we have
$\textstyle\quad\left\|\hat{\bm{T}}_{i}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\hat{\bm{T}}_{i}\right\|_{F}$
$=\left\|\left(\prod_{l=2}^{L}\bm{T}_{l}^{(i)}\right)(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\prod_{l=2}^{L}\bm{T}_{l}^{(i)}\right\|_{F}$
$\textstyle\leq\left\|\left(\prod_{l=2}^{L-1}\bm{T}_{l}^{(i)}\right)\left(\bm{T}_{L}^{(i)}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\bm{T}_{L}^{(i)}\right)\right\|_{F}$ (141)
$\quad\textstyle+\left\|\left(\prod_{l=2}^{L-2}\bm{T}_{l}^{(i)}\right)\left(\bm{T}_{L-1}^{(i)}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\bm{T}_{L-1}^{(i)}\right)\bm{T}_{L}^{(i)}\right\|_{F}+\cdots$
$\textstyle\quad+\left\|\left(\bm{T}_{2}^{(i)}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\bm{T}_{2}^{(i)}\right)\left(\prod_{l=3}^{L}\bm{T}_{l}^{(i)}\right)\right\|_{F}$
$\textstyle\leq\sum_{l=2}^{L}\sum_{j=1}^{2L-1}\left\|(\bm{\Sigma}_{1}^{(i)})^{j-1}\left(\bm{T}_{l}^{(i)}\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{1}^{(i)}\bm{T}_{l}^{(i)}\right)(\bm{\Sigma}_{1}^{(i)})^{2L-j-1}\right\|_{F}$
$\textstyle\leq\frac{\eta_{1}(2L-1)L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}\|\nabla G(\bm{W})\|_{F},$ (142)
where the last inequality follows from (35) and (65). Similarly, we have
$\left\|\hat{\bm{T}}_{i}\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{1}^{(i)}\hat{\bm{T}}_{i}\right\|_{F}\leq\frac{\eta_{1}L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}\|\nabla G(\bm{W})\|_{F}.$ (143)
Then we have
$\textstyle\quad\left\|\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\,\mathrm{BlkD}\left((\bm{\Sigma}_{1}^{(1)})^{L-1},\dots,(\bm{\Sigma}_{1}^{(p)})^{L-1},\bm{0}\right)\bm{\Psi}\hat{\bm{T}}\right\|_{F}$
$\textstyle=\left\|\mathrm{BlkD}\left(\hat{\bm{T}}_{1},\dots,\hat{\bm{T}}_{p},\bm{I}_{d_{1}-r_{\sigma}}\right)\left(\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}\right)\hat{\bm{T}}^{T}-\sqrt{\lambda}\,\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{0}\right)\bm{\Psi}\right\|_{F}$
$\textstyle\leq\left\|\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\,\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{0}\right)\bm{\Psi}\right\|_{F}$
$\textstyle\quad+\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}(\bm{\Sigma}_{1}^{(i)})^{2L-1}\hat{\bm{T}}_{i}^{T}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\right\|_{F}+\lambda\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}\bm{\Sigma}_{1}^{(i)}\hat{\bm{T}}_{i}^{T}-\bm{\Sigma}_{1}^{(i)}\right\|_{F}$
$\textstyle\leq\eta_{3}\left\|\nabla G(\bm{W})\right\|_{F}+\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\hat{\bm{T}}_{i}\right\|_{F}+\lambda\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{1}^{(i)}\hat{\bm{T}}_{i}\right\|_{F}$
$\textstyle\leq\left(\eta_{3}+\frac{p\eta_{1}(2L-1)L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}+\frac{\lambda p\eta_{1}L}{\sigma^{*}_{\min}}\right)\left\|\nabla G(\bm{W})\right\|_{F}=\eta_{4}\left\|\nabla G(\bm{W})\right\|_{F},$
where the first equality follows from the definition of $\bm{A}_{i}$ for each $i\in[p]$ in (54), the second inequality follows from (139), and the last inequality follows from (65), (142), and (143). ∎
B.4 Auxiliary Lemma for Proving 3
Lemma B.6.
Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary. Then there exists a constant $L_{G}>0$ such that
$\|\nabla G(\bm{W})\|_{F}\;\leq\;L_{G}\,\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}}),\quad\text{for all }\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\sigma_{\max}^{*},$
where
$L_{G}\;=\;\lambda L\;+\;2^{2L-1}L^{2}(\sigma_{\max}^{*})^{2L-2}\;+\;\sqrt{\lambda}\,y_{1}\,2^{L-2}L^{2}(\sigma_{\max}^{*})^{L-2}.$ (144)
Proof.
Suppose $\|\bm{W}-\bm{W}^{*}\|_{F}=\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})$ with $\bm{W}^{*}\in\mathcal{W}_{\bm{\sigma}^{*}}$, for each $l\in[L]$, we have
$\|\nabla_{l}G(\bm{W})\|_{F}$
$=$ $\|\nabla_{l}G(\bm{W})-\nabla_{l}G(\bm{W}^{*})\|_{F}$
$\leq$ $\lambda\|\bm{W}_{l}\!-\!\bm{W}^{*}_{l}\|_{F}\!+\!\|\bm{W}_{L:l+1}^{T}(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y})\bm{W}_{l-1:1}^{T}\!\!-\!\!\bm{W}_{L:l+1}^{*T}(\bm{W}_{L:1}^{*}\!\!-\!\!\sqrt{\lambda}\bm{Y})\bm{W}_{l-1:1}^{*T}\|_{F}$
$\leq$ $\lambda\|\bm{W}_{l}-\bm{W}^{*}_{l}\|_{F}+\|\bm{W}_{L:l+1}^{T}\bm{W}_{L:1}\bm{W}_{l-1:1}^{T}-\bm{W}_{L:l+1}^{*T}\bm{W}_{L:1}^{*}\bm{W}_{l-1:1}^{*T}\|_{F}$
$+\sqrt{\lambda}\|\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}-\bm{W}_{L:l+1}^{*T}\bm{Y}\bm{W}_{l-1:1}^{*T}\|_{F}.$ (145)
By 1, we have $\|\bm{W}_{l}^{*}\|\leq{\sigma}_{\max}^{*}\ \text{and}\ \|\bm{W}_{l}\|\leq\|\bm{W}_{l}^{*}\|+\|\bm{W}-\bm{W}^{*}\|_{F}\leq 2{\sigma}_{\max}^{*},\ \forall l\in[L].$ Using the triangle inequality and the matrix norm inequality $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|\|\bm{B}\|_{F}$, we have
$\left\|\bm{W}_{L:l+1}^{T}\bm{W}_{L:1}\bm{W}_{l-1:1}^{T}-\bm{W}_{L:l+1}^{*T}\bm{W}_{L:1}^{*}\bm{W}_{l-1:1}^{*T}\right\|_{F}$
$\leq$ $\left\|(\bm{W}_{l+1}-\bm{W}_{l+1}^{*})^{T}\bm{W}_{L:l+2}^{T}\bm{W}_{L:1}\bm{W}_{l-1:1}^{T}\right\|_{F}$
$+\left\|\bm{W}_{l+1}^{*T}(\bm{W}_{l+2}-\bm{W}_{l+2}^{*})^{T}\bm{W}_{L:l+3}^{T}\bm{W}_{l-1:1}^{T}\right\|_{F}+\cdots$
$+\left\|\bm{W}_{L:l+1}^{*T}\bm{W}_{L:1}^{*}\bm{W}_{l-2:1}^{*T}(\bm{W}_{l-1}-\bm{W}_{l-1}^{*})\right\|_{F}$
$\leq$ $L2^{2L-1}{\sigma_{\max}^{*}}^{2L-2}\|\bm{W}-\bm{W}^{*}\|_{F}.$ (146)
Using a similar telescope inequality, we have
$\sqrt{\lambda}\|\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}-\bm{W}_{L:l+1}^{*T}\bm{Y}\bm{W}_{l-1:1}^{*T}\|_{F}\leq\sqrt{\lambda}Ly_{1}2^{L-2}{\sigma}_{\max}^{*L-2}\|\bm{W}-\bm{W}^{*}\|_{F}.$ (147)
Plugging (146) and (147) into (145), we obtain
$\|\nabla G(\bm{W})\|_{F}$ $\leq\sum_{l=1}^{L}\|\nabla G_{l}(\bm{W})\|_{F}$
$\leq\left(\lambda L+2^{2L-1}L^{2}{\sigma}_{\max}^{*2L-2}+\sqrt{\lambda}y_{1}2^{L-2}L^{2}{\sigma}_{\max}^{*L-2}\right)\|\bm{W}-\bm{W}^{*}\|_{F}.$
Appendix C Linear Convergence under the Error Bound
Proposition 6 ( [23, 31, 37, 51]).
Suppose that 2 holds and the sequence $\{\bm{W}^{k}\}_{k\geq k_{1}}$ for some index $k_{1}\geq 0$ satisfies the following conditions:
  • (i)
    (Sufficient Decrease) There exists $\kappa_{1}>0$ with
    $F(\bm{W}^{k+1})-F(\bm{W}^{k})\leq-\kappa_{1}\|\bm{W}^{k+1}-\bm{W}^{k}\|_{F}^{2}.$
  • (ii)
    (Cost-to-Go Estimate) There exists $\kappa_{2}>0$ such that
    $F(\bm{W}^{k+1})-F(\bm{W}^{*})\leq\kappa_{2}\left(\mathrm{dist}^{2}(\bm{W}^{k},\mathcal{W})+\|\bm{W}^{k+1}-\bm{W}^{k}\|_{F}^{2}\right).$
  • (iii)
    (Safeguard) There exists a constant $\kappa_{3}>0$ such that $\|\nabla F(\bm{W}^{k})\|_{F}\leq\kappa_{3}\|\bm{W}^{k+1}-\bm{W}^{k}\|_{F}.$
Then, $\{F(\bm{W}^{k})\}_{k\geq k_{1}}$ converges Q-linearly to $F(\bm{W}^{*})$ and $\{\bm{W}^{k}\}_{k\geq k_{1}}$ converges R-linearly to $\bm{W}^{*}$ for some $\bm{W}^{*}\in\mathcal{W}$.
Most gradient-based algorithms generate sequences that satisfy the above supposition.
Appendix D Auxiliary Results
To bound the spectral gap between eigenvectors associated with repeated eigenvalues between two symmetric matrices, we introduce the Davis-Kahan theorem [46].
Lemma D.7.
Let $\bm{A},\hat{\bm{A}}\in\mathbb{R}^{p\times p}$ be symmetric with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{p}$ and $\hat{\lambda}_{1}\geq\cdots\geq\hat{\lambda}_{p}$ respectively. Fix $1\leq r\leq s\leq p$ and assume that $\min\{\lambda_{r-1}-\lambda_{r},\lambda_{s}-\lambda_{s+1}\}>0$, where $\lambda_{0}:=\infty$ and $\lambda_{p+1}:=$-$\infty$. Let d:=sr+1d:=s-r+1 and $\bm{V}=(\bm{v}_{r},\bm{v}_{r+1},\dots,\bm{v}_{s})\in\mathbb{R}^{p\times d}$ and $\hat{\bm{V}}=(\hat{\bm{v}}_{r},\hat{\bm{v}}_{r+1},\dots,\hat{\bm{v}}_{s})\in\mathbb{R}^{p\times d}$ have orthonormal columns satisfying $\bm{A}\bm{v}_{j}=\lambda_{j}\bm{v}_{j}$ and $\hat{\bm{A}}\hat{\bm{v}}_{j}=\hat{\lambda}_{j}\hat{\bm{v}}_{j}$ for all $j=r,r+1,\ldots,s$. Then, there exists an orthogonal matrix $\bm{O}\in\mathcal{O}^{d}$ such that
$\|\hat{\bm{V}}\bm{O}-\bm{V}\|_{F}\leq\frac{4\|\hat{\bm{A}}-\bm{A}\|_{F}}{\min\{\lambda_{r-1}-\lambda_{r},\lambda_{s}-\lambda_{s+1}\}}.$
Lemma D.8 (Mirsky’s Inequality [38]).
Let $\bm{X},\tilde{\bm{X}}\in\mathbb{R}^{m\times n}$ be arbitrary matrices with singular values
$\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{l}\quad\text{and}\quad\tilde{\sigma}_{1}\geq\tilde{\sigma}_{2}\geq\cdots\geq\tilde{\sigma}_{l},$
respectively, where $l=\min\{m,n\}$. Then, for any unitarily invariant norm (e.g., the Frobenius norm), we have
$\left\|\mathrm{diag}(\tilde{\sigma}_{1}-\sigma_{1},\dots,\tilde{\sigma}_{l}-\sigma_{l})\right\|\leq\|\tilde{\bm{X}}-\bm{X}\|.$